### v1

# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

Abstract
Batch normalization (BN) is a ubiquitous operation in deep neural networks used primarily to achieve stability and regularization during network training. BN involves feature map centering and scaling using sample means and variances, respectively. Since these statistics are being estimated across the feature maps within a batch, this problem is ideally suited for the application of Stein’s shrinkage estimation, which leads to a better, in the mean-squared-error sense, estimate of the mean and variance of the batch. In this paper, we prove that the Stein shrinkage estimator for the mean and variance dominates over the sample mean and variance estimators in the presence of adversarial attacks when modeling these attacks using sub-Gaussian distributions. This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and use it in image classification (segmentation) tasks with and without adversarial attacks. We present SOTA performance results using this Stein corrected batch norm in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
1 Introduction
Deep Neural Networks are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and sciences. In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of deep neural networks. Deep network training is well known to be time intensive and unstable. Batch normalization (BN) is a particularly useful tool in achieving numerical stability and regularization during network training. BN typically involves centering and scaling the feature maps. The centering and scaling operations involve computing the sample means and variances, respectively, from the samples/members in the batch. This task is ideally suited for the application of Stein’s shrinkage method, which is well known in the field of statistics. Before delving into the Stein’s shrinkage method, let us briefly review the widely popular batch normalization in deep learning.
Batch normalization was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift Ioffe-ICML15 . Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) ba2016layernormalization and Transformers, spectral normalization in generative adversarial networks (GANs) Miyato-ICLR18 , and many others. Improvements over BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training has also been proposed in Zhang_2021_CVPR and batch normalization preconditioning in Lange-JMLR2022 , wherein the pre-conditioner is constructed using batch statistics. These normalization algorithms do not address the effectiveness of improvements in stability or convergence of training in the presence of adversarial attacks.
In this paper, we address the problem of developing a BN that leads to better prediction accuracy from the deep network used in classification in the presence of adversarial attacks. Adversarial attacks have been modeled via additive bounded perturbations MauryaArxiv22 . Some of the popular adversarial attack methods that are popular and satisfy the bounded perturbation condition are FGSM (fast gradient sign method) szegedy2014 and an iterative variaint of PGD (projected gradient descent). Numerous articles exist in the literature that depict ways to overcome a drop in the accuracy of classifiers in the presence of adversarial attacks. Our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, however, to adapt an existing shrinkage estimator specifically, the well known James-Stein shrinkage estimator and prove that it retains its dominance over the sample mean and variance in the presence of adversarial attacks. Dominance here refers to smaller mean squared error (MSE) – between the estimate and the true value – when compared to the sample mean and variance. This smaller MSE implies a better BN in each layer and as we demonstrate empirically, a better stability and prediction accuracy from the deep network. By making BN stable in the presence of adversarial attacks this work could have potential applications that might have societal impact.
1.1 A Brief on the James-Stein Shrinkage Estimator
In the following we present a brief introduction to the James-Stein (JS) shrinkage estimator and motivate it’s aptness in the context of batch normalization in deep neural networks. This will provide the necessary background and motivation for our work presented in this paper.
Shrinkage estimation is a parameter estimation technique that is primarily used for variance reduction and to improve parameter estimation accuracy by shrinking the parameter towards a fixed target. For instance, shrinking the estimate of the mean of a Gaussian toward zero. It is especially useful in high-dimensional and scarce data settings. The shrinkage, of course, leads to a biased estimate, but with smaller variance, which is sought after in many applications. The James-Stein (JS) shrinkage estimator was proposed in the field of statistics as an estimator of the mean for samples drawn from a Gaussian distribution and has been shown to dominate the maximum likelihood estimator (MLE) in terms of the risk (mean-squared error) James-Stein61 . Dominance here refers to the risk associated with the JS-estimator being at most as much as that of the MLE and that there exists at least one set of parameter values for which it is strictly less. More precisely, if $\mathbf{X}$ is a $p$=dimensional random vector whose mean is to be estimated and it is assumed that $\mathbf{X}\sim N(\mathbf{\mu},\mathbf{I})$, then, $E(||\mathbf{\mu}^{JS}-\mathbf{\mu}||^{2})\leq E(||\mathbf{\mu}^{MLE}-\mathbf{\mu}||^{2})$. It is important to note that the JS-estimator improves the overall risk of estimating the mean (parameter) vector and not the risk of the individual elements of the mean vector.
At the foundation of the JS-shrinkage estimator lies what is known as Stein’s (identity) lemma, which characterizes the normal distribution. Stein’s (identity) lemma is also at the foundation of the Stein’s method which is used for distributional comparisons. We will not discuss this latter well-known method in statistics here and instead refer the interested reader to a recent comprehensive survey Anastasiou23 .
Since its inception, the JS-shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family Brown66 , exponential family Hudson78AOS , Gamma distribution DasGupta86AOS , distributions with quadratic varaince function XianchaoXie-Brown-AOS16 and FathiAOS2022 for relaxing the Gaussian assumption in high dimensions. For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to Kubokawa24 . In addition, more recently, Stein’s shrinkage estimate has been generalized to distributions on manifold-valued random variables Yang-Doss-Vemuri22 .
1.2 Relevance of JS-estimator to BN
Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to batch normalization in deep neural networks and how it can be applied to improve the stability and accuracy of predictions. A wide range of normalization techniques have been developed to facilitate the training of deep neural networks by normalizing intermediate feature distributions and improving the stochastic optimization dynamics. One of the earliest and most influential approaches is BN, introduced in Ioffe-ICML15 , which normalizes features using statistics computed over the entire mini-batch. This method was originally motivated by the desire to reduce internal covariate shift, though subsequent work has highlighted its strong impact on the optimization landscape conf/nips/SanturkarTIM18 as a key factor behind its effectiveness. In the following, we give a brief and precise description of different types of normalization that are in vogue in literature.
Let $X\in\mathbb{R}^{N\times C\times H\times W}$ denote the feature tensor at a given layer of a deep neural network, where $N$ – batch size, $C$ – number of channels (feature maps), $H,W$ – spatial dimension. Formally, many normalization techniques follow a common template. Each element $x_{i}\in X$ is normalized as: $\hat{x_{i}}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}}$, where $i=(i_{N},i_{C},i_{H},i_{W})$ is a multi-index ranging over the tensor dimensions and $\mu_{i},\sigma^{2}_{i}$ are the empirical mean and variance computed over a subset $S_{i}\subseteq X:$ $\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\ \ \sigma_{i}^{2}=\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2}.$ The choice of $S_{i}$ determines the normalization scheme and the available choices are the following: (i) Batch Normalization (BN)  Ioffe-ICML15 : $S_{i}^{BN}={(n^{\prime},h^{\prime},w^{\prime})|n^{\prime}\in[1,N],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed over all samples in the batch and all spatial positions, but separately for each channel $i_{c}.$ (ii) Layer Normalization (LN)  ba2016layernormalization : $S_{i}^{LN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in[1,C],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed independently for each sample (fixed $i_{N}$) across all channels and spatial locations. (iii) Instance Normalization (LN)  ulyanov2017instancenormalizationmissingingredient : $S_{i}^{IN}={(h^{\prime},w^{\prime})|h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Normalization is per-sample and per-channel — only over the spatial dimensions. (iv) Group Normalization (GN)  wu2018groupnormalization : $S_{i}^{GN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in C_{\text{group}},h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Channels are split into groups; statistics are computed over spatial dimensions and within the channel group. After normalization, a learnable affine transformation is applied to preserve representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
While normalization methods like LayerNorm, InstanceNorm, and GroupNorm have been introduced to mitigate batch-size dependence, BatchNorm remains a central component in many modern architectures and continues to attract interest both in theory and practice. A key observation is that BatchNorm estimates a large number of mean and variance values simultaneously — one pair for each channel. This can be naturally viewed as a multitask estimation problem: we are estimating a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$ each based on a finite number of observations.
In such settings, James–Stein shrinkage estimators offer a statistically grounded way to improve estimation by borrowing strength across tasks — in this case, across channels. The classical result suggests that when estimating multiple parameters jointly under squared loss, shrinking individual estimates toward their grand mean can reduce overall error. Applying this idea in the context of BatchNorm could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
Having established adequate motivation for application of Stein’s shrinkage to BN in a deep network, we are now ready to narrate the rest of the story. But before that, a word on what is to follow. Section 2 delves into the presentation of the appropriate shrinkage formulas for the BN. This is followed by the presentation of theoretical results on the dominance of Stein’s shrinkage formulas in the presence of adversarial attacks in Section 3. In Section 4, we present experimental results that compare the performance of the proposed shrinkage to SOTA on several benchmark data sets. Finally in Section 5, we present the discussion and conclusions.
2 Stein’s Shrinkage Applied to Batch Normalization
In this section, we apply the Stein’s shrinkage to the BN and present the formulas for calculation of the batch normalization in a deep neural network. We will limit ourselves primarily to convolutional neural networks (CNNs) but the formulas can be applied to any architecture with some minor changes.
Let $X\sim\mathcal{N}_{p}(\theta,I_{p})$, where $\theta\in\mathbb{R}^{p}$ is an unknown parameter vector and $I_{p}$ is the $p$-dimensional identity matrix. The James Stein shrinkage estimator of the unknown parameter $\theta$ is given by, $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X$. It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator but is obviously a biased estimator. However, a smaller mean squared error is desirable in many applications. Most recently, JS-estimator for the calculation of the mean in BN was used in khoshsirat2023improvingnormalizationjamessteinestimator . However, the authors in khoshsirat2023improvingnormalizationjamessteinestimator used the same JS-estimator formula to shrink the mean as well as the variance in BN. Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the sample variance is not normally distributed but instead follows a scaled chi-squared distribution - a special case of the Gamma distribution. To address this, we adopt a principled alternative.
More specifically, for mean correction, we adopt the James-Stein shrinkage under the assumption that individual elements of the feature tensor in the BN are normally distributed and homoscedastic. The homoscedasicity assumption simplifies our calculations and we will address relaxing this assumption in future. Under this assumption, the sample mean is normally distributed, and the sample variance follows a chi-squared distribution (i.e., a Gamma distribution): $\hat{\sigma}_{i}^{2}\sim\frac{\sigma_{i}^{2}}{n}\cdot\chi^{2}_{n-1}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$, which enables us to apply shrinkage to the scale parameter of the Gamma distribution DasGupta86AOS . Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ independently for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$. Then JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right]$ (1)
To incorporate the James-Stein correction into Batch Normalization, we compute the statistics for each channel and apply the shrinkage formulas as outlined below. Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions given by:
$\mu_{C}=\frac{1}{NHW}\sum_{n,h,w}x_{n,c,h,w},\quad\sigma^{2}_{C}=\frac{1}{NHW}\sum_{n,h,w}(x_{n,c,h,w}-\mu_{C})^{2}.$
We then apply shrinkage to these estimates. The James-Stein corrected mean and variance are:
Stein Shrinkage Forumals for Mean & Variance $\mu^{\mathrm{JS}}_{C}$ $=\left(1-\frac{(c-2)\sigma^{2}_{\mu_{C}}}{\|\mu_{C}\|^{2}_{2}}\right)\mu_{C},\quad{\sigma_{C}^{2}}^{\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+\tilde{c}\cdot V,$ (2) $\text{where }V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad\tilde{c}\in\left[0,\dfrac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$
Here, $\sigma^{2}_{\mu_{C}}$ denotes the estimated variance of the corresponding vector $\mu_{C}$.
Finally, the normalized output is given by:
Final Normalized (scaled and shifted) Output $\hat{x}=\frac{x-\mu^{\mathrm{JS}}}{\sqrt{{\sigma^{2}}^{\mathrm{JS}}+\epsilon}},\quad y=\gamma\hat{x}+\beta.$ (3)
Equations 2 and 3 respectively are the formulas we advocate for use in BN corrections and will be used in all our experiments. We are now ready to present the theoretical results of dominance of these shrinkage estimators in the presence of adverserial attacks.
3 Dominance of Stein Shrinkage in the Presence of Adverserial Attacks
In this section, we present the main theoretical contribution of our work: the James–Stein estimator continues to exhibit dominance even under adversarial perturbations. We begin by formalizing how adversarial attacks can be modeled probabilistically. Then, we introduce the necessary preliminaries and outline the main ideas behind the proof.
We model adversarial perturbations using a sub-Gaussian distribution, which is well-motivated both practically and theoretically. Adversarial perturbations are typically bounded in magnitude (e.g., under an $l_{p}$-norm constraint) to ensure the model remains learnable — a standard assumption in the literature cohen2019certifiedadversarialrobustnessrandomized ; zhai2019adversariallyrobustgeneralizationjust . Since bounded random variables are inherently sub-Gaussian, this provides a natural modeling choice. Additionally, the sub-Gaussian model enables a tractable analysis via the use of concentration inequalities and allows one to work with a well-defined variance proxy. Following prior workMauryaArxiv22 , we assume that each adversarial perturbation vector has coordinates bounded in $[-\varepsilon,\varepsilon]$ and can be represented as a zero-mean sub-Gaussian vector with variance proxy $2\varepsilon^{2}$: $Y\sim\text{subG}(2\varepsilon^{2}),\quad\text{i.e., for all }v\in\mathbb{R}^{p}\text{ with }\|v\|=1,\ \ \text{and}\ \ \mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\frac{2\varepsilon^{2}\lambda^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$ This assumption allows us to model the observed vector as the sum of noise-free data and adversarial noise: $Z=X+Y,\ \ \ X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$ For simplicity we assume that $X$ and $Y$ are independent, which is currently a limitation of our model. Other noise models that are data dependent will be considered in our future work.
3.1 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Normally Distributed Random Variables
Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\in\mathbb{R}^{p}$ is independent, mean-zero sub-Gaussian vector with a known proxy variance $2\varepsilon^{2}.$ Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{\mathrm{MLE}}$ and the JS-estimator
$\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{(p-2)\hat{\sigma}^{2}}{\|Z\|^{2}_{2}}\right)Z.$
under squared error loss. It is well-known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{\mathrm{MLE}}$ when $Y=0$ and $p\geq 3$ James-Stein61 . We aim to show that this dominance remains valid even under additive sub-Gaussian noise, provided the proxy variance $2\varepsilon^{2}$ is sufficiently small.
We first observe that it suffices to consider the case $\sigma^{2}=1$ The result for general $\sigma^{2}$ then follows by a simple scaling argument. Our first key result is the following dominance theorem for the Gaussian random variable:
JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gaussian Model
Theorem 3.1.
Let $\theta\in\mathbb{R}^{p}$ be an unknown parameter vector. Suppose we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent, mean-zero sub-Gaussian vector with parameter $2\varepsilon^{2}$, i.e., for all unit vectors $v\in\mathbb{R}^{p}$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\varepsilon^{2}\lambda^{2}\right)$. Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the maximum likelihood estimator by $\hat{\theta}^{\mathrm{MLE}}=Z$. Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the MLE in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{MLE}},\theta).$
We now provide a sketch of the proof and refer the reader to the Appendix for more details.
Sketch of Proof.
We can rewrite the risk of the JS-estimator as:
$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
The last two terms in equation 4 can be rewritten as:
$-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\right)$
Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that: $\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$ Hence, the risk of the James–Stein estimator is strictly less than that of the MLE for $p\geq 3$ completing the proof. ∎
We are now ready to address the issue of shrinkage estimation of the variance parameter in the above normally distributed random variable with additive sub-Gaussian perturbation.
3.2 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Gamma Distributed Random Variables
In many applications, empirical variances follow a scaled chi-squared distribution, which is a special case of the gamma distribution. Given observations $x_{ij}$, define the empirical variance: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$. Thus, estimating $\sigma_{i}^{2}$ reduces to estimating the scale parameter $\beta_{i}$ in a gamma distribution with fixed shape parameter $\alpha$. For this task, we can use the following theorem from literature:
Theorem 3.2 (JS-Estimator Shrinkage Dominance for the Gamma Model DasGupta86AOS ).
[…]
Now, applying the above theorem to empirical variances, we obtain: $\hat{\sigma}_{i}^{2,\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+cV,\quad V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad c\in\left[0,\frac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$ We now present our second key result namely, that this dominance persists under additive sub-Gaussian noise for the Gamma distributed random variable.
JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gamma Model
Theorem 3.3.
[…]
Remark 3.4.
The admissible range for $c$ under sub-Gaussian perturbations includes the classical interval, so standard shrinkage remains valid even in noisy settings.
Sketch of Proof.
[…]
We begin by expressing the risk of $\hat{\beta}^{\,\mathrm{JS}}$ as:
[…]
Thus, the difference in risks depends on the sign of the second term in the above expression expressed as: $c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$ To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term. We consider two main cases:
  • Case 1:
    $c<0$ and
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (5)
  • Case 2:
    $c>0$ and
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (6)
In both cases outlined above, the proof fundamentally relies on an application of Stein’s lemma for the Gamma distribution given below. Since each $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ with known shape and rate parameters, we can invoke the Stein’s identity given in the following lemma:
Lemma 3.5 (Stein’s Lemma for the Gamma Distribution).
Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty$. Then the following (Stein’s) identity holds: $\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)]$.
We apply Stein’s identity to the function $h(x)=V$, where $V$ is the geometric mean defined by $V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}$. Using the Stein’s identity for the Gamma distributed random variable, which is given by, $\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$ we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get: $\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality. These tools allow us to bound the expressions for a given choice of $c$ and derive the dominance for the JS-estimator of the variance parameter of the Normally distributed random variable under sub-Gaussian perturbations.
Further technical details and precise bounds are deferred to the appendix. ∎
4 Experiments
In this section, we present several experimental results demonstrating the performance of BN with our proposed shrinkage estimators in both image classification and segmentation tasks, under standard and adversarial settings. To simulate adversarial attacks, we added noise samples drawn from a sub-Gaussian distribution with a fixed variance. The sub-Gaussian density is given by $f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$ derived from a mixture of Lévy and Gaussian distributions, and the samples were drawn using the inverse transform sampling technique. GEORGIOU20063061 . For classification, we employed the ResNet-9 he2015deepresiduallearningimage architecture on the CIFAR-10  cifar10 dataset, as well as a 3D CNN model on the PPMI dataset ppmi-data . For object segmentation, we used the HRNetV2 wang2020deep model applied to the Cityscapes datasetCordts2015Cvprw . Since BN is known to become unstable when the batch size is small, we report results for varying batch sizes and sub-Gaussian noise levels. Our goal is to demonstrate that our shrinkage-based method is more robust to adversarial perturbations and offers greater overall stability. We remind the reader that our goal here is not to improve the SOTA results on these data in the literature, but to demonstrate the performance gains from the use of BN in the presence (absence) of adversarial attacks with the correct use of Stein shrinkage compared to vanilla BN and other forms of shrinkage applied to BN. Throughout all the experiments, we avoided training-time tricks and focused on obtaining a honest comparison.
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
Standard (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
Standard (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
Our (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
Lasso (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
Lasso (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
Lasso corrected (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
Ridge (32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
Ridge (64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
Ridge (128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
Table 1: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
In the experiments, we compare our approach mainly with that of Khoshsirat et al. khoshsirat2023improvingnormalizationjamessteinestimator , aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes. In addition, with the correct variance shrinkage, one can achieve higher accuracy in the presence of adversarial perturbations. We also tested a variant using only the mean shrinkage estimator and no shrinkage on variance. Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are commonly used shrinkage methods. Detailed formulas for these estimators and complete experimental setup are provided in the Appendix B. Note that in all the tables depicting the results, we have highlighted the best performing BN method (ours) for small batch sizes that are preferred from a computational efficiency perspective as well as when data are scarce.
4.1 Application to Cifar10 Data
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. The results in Table 1 confirm that our proposed BN significantly outperforms the baselines in noisy settings across all batch sizes. Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in khoshsirat2023improvingnormalizationjamessteinestimator , our approach shows consistent improvement due to the correct development and use of variance shrinkage. The method with only mean correction (see Appendix) performs worse than our full model, highlighting the importance of variance correction. Finally, Lasso and Ridge-based approaches do not improve noise performance, showing behavior similar to the standard BN.
4.2 Application to Cityscapes Data
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
Standard (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
Our (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
Lasso (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
Lasso (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
Ridge (8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
Ridge (16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset—a more challenging task involving high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. As in the previous experiment, we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. This setup is particularly relevant for segmentation, where maintaining accurate local feature distributions is crucial for pixel-level predictions. As shown in Table 2, standard BN fails to retain meaningful performance even under mild noise (e.g., below $10$ mIoU score at $3\%$ corruption). In contrast, our proposed method and its mean corrected variant (see the appendix) significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings. Regularization-based approaches (Lasso and Ridge) show negligible improvement under noise and have a behavior similar to that of standard BN.
4.3 Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN method on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline (see Appendix B) using the FSL software FSLpack . DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local Gaussian fit. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization FORTIN2017149 is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used. The data set is highly imbalanced and contains FA maps from 360 Control and 130 PD brain scans. This imbalance is addressed using mixup-basedDBLP:journals/corr/abs-1710-09412 augmentation with sampling bias favoring the smaller sized class (PD). Using the same setup as in previous experiments and a 5-fold cross-validation to compare 3D CNN with different variants of 3D BN, we get results summarized in the table 3. These results further support our earlier observations, even in this challenging classification task.
BN Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge (10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge (16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on PPMI data. Values depict mean accuracies $\pm$ standard deviation.
5 Conclusions
In this paper, we presented a novel theory involving the proof of dominance of the Stein’s shrinkage estimators of mean and variance of a Gaussian random variable in the presence of adversarial attacks. We applied Stein’s shirnkage to BN in a variety of deep networks and demonstrated superior performance, in comparison to other BN methods, on three benchmark data sets. One possible explanation for the improved performance is the connection between model performance under attack and the Lipschitz properties of the network DBLP:journals/corr/abs-2003-02460 . Since BN is known to affect the Lipschitz constant conf/nips/SanturkarTIM18 , it would be interesting to investigate the influence of our BN method on this aspect in future work.
6 Acknowledgments
This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099. (Corresponding author: Baba C. Vemuri.)
Appendix A Technical Appendices and Supplementary Material
In this Appendix, we present the detailed proofs of the two key theorems 3.1 and 3.3 and additional experiments showing Stein’s correction applied only to the mean and not the variance. These experimental results are depicted here in this Appendix as additional rows added in to the three tables appearing in the main body of the paper. Also included are further experimental details.
Theorem 3.1 Proof:.
We denote MLE estimator as $\hat{\theta}^{\mathrm{MLE}}=z=\left(z_{1},\ldots z_{p}\right).$
Then James - Stein Estimator is $\hat{\theta}^{\mathrm{JS}}=z-\frac{(p-2)z}{z^{\top}z}=z\left(1-\frac{p-2}{z^{\top}z}\right)=\hat{\theta}^{\mathrm{MLE}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{MLE}\top}\hat{\theta}^{\mathrm{MLE}}}\right).$
The risk of James - Stein Estimator equals to
$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$
$=\mathbb{E}\left(\left\{(z-\theta)-\frac{(p-2)z}{z^{\top}z}\right\}^{\top}\left\{(z-\theta)-\frac{(p-2)z}{z^{\top}z}\right\}\right)=$
$=\mathbb{E}\left((z-\theta)^{\top}(z-\theta)\right)-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}{\mathbb{E}\left(\frac{1}{z^{\top}z}\right)}=$
$=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)$
Let us consider the last two terms:
$-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
$=-2(p-2)\left(1-\mathbb{E}\left(\frac{z^{\top}\theta}{z^{\top}z}\right)\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
$=(p-2)\left(-2+\mathbb{E}\left(\frac{2z^{\top}\theta}{z^{\top}z}\right)+\mathbb{E}\left(\frac{(p-2)}{z^{\top}z}\right)\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\right)$
We want to show that $\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)<2$
Let us denote $\tilde{x}=2z^{\top}\theta+p-2\quad$ and $\tilde{y}=\frac{1}{z^{\top}z}.$
By applying Cauchy–Schwarz inequality we get,
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)=\mathbb{E}(\tilde{x}\cdot\tilde{y})\leq|\mathbb{E}(\tilde{x}\cdot\tilde{y})|\leq\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}.$
Let us calculate each term separately:
$\mathbb{E}\left(\tilde{x}^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta+p-2\right)^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta\right)^{2}+4\left(z^{\top}\theta\right)(p-2)+(p-2)^{2}\right)=$
$=\mathbb{E}\left(4\left(x^{\top}\theta+y^{\top}\theta\right)^{2}+4\left(x^{\top}\theta+y^{\top}\theta\right)(p-2)\right)+(p-2)^{2}=$
$=4\mathbb{E}\left((x^{\top}\theta)^{2}\right)+8\mathbb{E}\left((x^{\top}\theta)(y^{\top}\theta)\right)+4\mathbb{E}\left((y^{\top}\theta)^{2}\right)+4(p-2)\mathbb{E}\left(x^{\top}\theta\right)+(p-2)^{2}=$
$=4Var\left(x^{\top}\theta\right)+4\left(\mathbb{E}\left(x^{\top}\theta\right)\right)^{2}+4Var\left(y^{\top}\theta\right)+4\left(\mathbb{E}\left(y^{\top}\theta\right)\right)^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=$
$=4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+4Var\left(y^{\top}\theta\right)+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}\leq$
$4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+8\epsilon^{2}||\theta||_{2}^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}.$
So, $\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\leq\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}.$
Note that,
$\mathbb{E}\left(z^{\top}z\right)=\mathbb{E}\left(\sum_{i}z_{i}^{2}\right)=\sum_{i}\mathbb{E}\left(z_{i}^{2}\right)=\sum_{i}Var(z_{i})+\sum_{i}\left(\mathbb{E}(z_{i})\right)^{2}\geq p+||\theta||_{2}^{2}.$
Using Jensen inequality, firstly for $f(x)=\frac{1}{x}$ and then for $g(x)=x^{2}$ we get
$\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}=\sqrt{\mathbb{E}\left(\frac{1}{(z^{\top}z)^{2}}\right)}\leq\sqrt{\frac{1}{\mathbb{E}\left((z^{\top}z)^{2}\right)}}\leq\frac{1}{\mathbb{E}\left(z^{\top}z\right)}\leq\frac{1}{p+||\theta||_{2}^{2}}.$
So, overall we get
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\leq\frac{\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}}{p+||\theta||_{2}^{2}}\leq 2.$
The final inequality can be established by direct computation. ∎
Theorem 3.3 Proof:.
[…]
Note that,
[…]
So, we need to show:
$c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right)<0.$
We consider two cases depending on the sign of the constant $c$:
  • Case 1:
    $c<0$ and
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (7)
  • Case 2:
    $c>0$ and
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (8)
Let us derive some facts that will be helpful in the proof.
Firstly, we aim to show that for each $i$, the following inequality holds:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta.$
Note that:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)=1-\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}<0\right).$
Let us define:
$W_{i}:=\frac{2}{n}\sum_{j=1}^{n}(X_{ij}-\bar{X}_{i})(Y_{ij}-\bar{Y}_{i})=\frac{2}{n}\sum_{j=1}^{n}w_{j}.$
The centered random variables $X_{ij}-\bar{X}_{i}$ and $Y_{ij}-\bar{Y}_{i}$ are sub-Gaussian with parameters
$\sigma_{x,i}^{2}\left(1+\frac{1}{n}\right)\quad\text{and}\quad\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right),$
respectively. Their product $w_{j}$ is a sub-exponential random variable with parameters
$\nu^{2}=C_{1}\cdot\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2},\quad b=C_{2}\cdot\sqrt{\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2}}.$
Reminder. A random variable $Z$ is sub-exponential if:
$\mathbb{E}[e^{\lambda Z}]\leq\exp\left(\frac{\lambda^{2}\nu^{2}}{2}\right),\quad|\lambda|<\frac{1}{b}.$
Bernstein Inequality for Sub-Exponential Random Variables. Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential variables with parameters $(\nu,b)$. Then for any $t>0$:
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\leq-t\right)\leq\begin{cases}\exp\left(-\dfrac{nt^{2}}{2\nu^{2}}\right),&\text{if }0\leq t\leq\dfrac{\nu^{2}}{b},\\ \exp\left(-\dfrac{nt}{2b}\right),&\text{if }t>\dfrac{\nu^{2}}{b}.\end{cases}$
Applying this inequality to $\frac{1}{n}\sum_{j}w_{j}$, we obtain:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}<0\right)=\mathbb{P}\left(\frac{2}{n}\sum_{j}w_{j}<-\hat{\sigma}^{2}_{y,i}\right)=\mathbb{P}\left(\frac{1}{n}\sum_{j}w_{j}<-\frac{\hat{\sigma}^{2}_{y,i}}{2}\right),$
which can be bounded by:
$\leq\begin{cases}\exp\left(-\dfrac{n(\hat{\sigma}^{2}_{y,i})^{2}}{8C_{1}\sigma_{x,i}^{2}\sigma_{y,i}^{2}(1+\frac{1}{n})^{2}}\right),&\text{if }\frac{\hat{\sigma}^{2}_{y,i}}{2}\leq\dfrac{\nu^{2}}{b},\\ \exp\left(-\dfrac{n\hat{\sigma}^{2}_{y,i}}{4C_{2}\sqrt{\sigma_{x,i}^{2}\sigma_{y,i}^{2}(1+\frac{1}{n})^{2}}}\right),&\text{otherwise}.\end{cases}$
Therefore, we conclude:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta,$
where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$. Thus, the desired non-negativity condition holds with high probability in the large sample regime. We will use this fact in the proof of case 2. Now we are ready to consider each case separately.
Case 1. We examine the first term in more detail:
$2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]=2\sum_{i=1}^{p}\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}-\beta_{i}\right)V\right]$ (9)
$=2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)$
Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can apply Stein’s lemma for the Gamma distribution:
Lemma A.1 (Stein’s Lemma for the Gamma Distribution).
Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that
$\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty.$
Then the following identity holds:
$\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)].$
We apply Stein’s identity to the function $h(x)=V$, where
$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}.$
Using the Gamma Stein identity:
$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$
we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ as follows:
$\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}.$
Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
In particular, we obtain the inequality:
$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$
Using this bound equation (9) becomes:
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}V\right]-\beta_{i}\cdot\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)=$
$=2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$
Let’s denote
$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$
Overall, we get lower bound for :
$2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\mathbb{E}[V^{2}]\geq-\frac{2p}{\alpha(\alpha+1)}\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]+pc\mathbb{E}[V^{2}]\geq$
$\geq-p\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]\left(\frac{2}{\alpha(\alpha+1)}+c\right)>0$
Last inequality appears due to assumption. So we have:
$c<0\ \text{and}\ \frac{2}{\alpha(\alpha+1)}+c<0\Rightarrow c\leq-\frac{2}{\alpha(\alpha+1)}.$
Case 2. We assume $c>0$ and we want to show
$02\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<$ 0 .
Recall that using Bernstein’s inequality we get with high probability, the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative. So, combining this with Stein’s lemma for $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ we get:
$\mathbb{E}(V)$ $=\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}\left(V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}\right)$
$\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}(V)$
$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$
$\Rightarrow\quad-\mathbb{E}(V)$ $\leq-\frac{1}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\cdot\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V).$
Plug in this bound and get inequality for  (9):
$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)$ $=2\sum_{i=1}^{p}\mathbb{E}\left(\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right)\right)\right)$
$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$
Using our previous notation $\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2},$
$=\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$
Overall for (8) we get:
$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})\leq$
$\leq\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)+pc\,\mathbb{E}(V^{2})<0$
The last inequality we want to establish by finding the appropriate $c$. So, we will rewrite it in a more suitable way:
$c\,\mathbb{E}(V^{2})<\frac{2}{\alpha p+1}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)-\frac{2}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)$
$c<\frac{2}{\alpha p+1}\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}-\frac{2}{\alpha+1}\cdot\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}$ (10)
Let’s bound each term independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, then we bound second term
$\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)\geq\mathbb{E}(V^{2})\quad\Rightarrow\quad-\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}\leq-1.$ (11)
Also, using Cauchy–Schwarz inequality:
$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$
For the first term we get:
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}\cdot\sqrt{\mathbb{E}(V^{2})}}{\mathbb{E}(V^{2})}=\sqrt{\frac{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}{\mathbb{E}(V^{2})}}$ (12)
By applying Jensen’s inequality for $f(x)=\sqrt{x}$ – concave, we have:
$\sqrt{\mathbb{E}(V^{2})}\geq\mathbb{E}(|V|)=\mathbb{E}(V)\Rightarrow\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}$ (13)
Now,
$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq$
Again, using Jensen’s inequality for $f(x)=\exp{x}$
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$
We know that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
$\mathbb{E}(\log\sigma_{x,i}^{2})\ \geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$
Plug in this inequality to the derivations above we get:
$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$
For (13) we get:
$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$
Also, we can directly calculate numerator of (12):
Firstly,
$\mathbb{E}(\hat{\sigma}_{x,i}^{4})=Var(\hat{\sigma}_{x,i}^{2})+(\mathbb{E}(\hat{\sigma}_{x,i}^{2}))^{2}=\alpha\beta_{i}^{2}+\alpha^{2}\beta_{i}^{2}=\alpha\beta_{i}^{2}(1+\alpha).$
$\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2})=\alpha^{2}\beta_{i}\beta_{j}.$
$\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}=\sqrt{\sum_{i=1}^{p}\mathbb{E}(\hat{\sigma}_{x,i}^{4})+2\sum_{1\leq i<j\leq p}\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2}})=$
$=\sqrt{\sum_{i=1}^{p}\alpha\beta_{i}^{2}(1+\alpha)+2\sum_{1\leq i<j\leq p}\alpha^{2}\beta_{i}\beta_{j}}=\sqrt{\alpha(1+\alpha)\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}\left((\sum_{i=1}^{p}\beta_{i})^{2}-\sum_{i=1}^{p}\beta_{i}^{2}\right)}=$
$=\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$
Therefor, for (12) we get upper bound:
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$
Using this inequality and (11) finally for (10) we get:
$c<\frac{2}{\alpha p+1}\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}-\frac{2}{\alpha+1}=$
$=\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$
Note, that
$\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}\geq\frac{p}{\sum_{i=1}^{p}\frac{1}{\beta_{i}}}\Rightarrow\frac{1}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\leq\frac{\sum_{i=1}^{p}\frac{1}{\beta_{i}}}{p}.$
Also, by Jensen’s inequality we get
$\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\beta_{i}}\leq\frac{1}{\frac{1}{p}\sum_{i=1}^{p}\beta_{i}}=\frac{p}{\sum_{i=1}^{p}\beta_{i}}.$
Hence,
$\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\leq\frac{1}{p}\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}\left(\sum_{i=1}^{p}\frac{1}{\beta_{i}}\right)$
$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$
So, dominance holds for
$0<c<\frac{2p}{(\alpha p+1)}\exp{\left(\frac{1}{\alpha}\right)}\sqrt{\frac{1}{\alpha}+1}-\frac{2}{\alpha+1}.$
Note that under regular conditions Stein’s estimate dominance holds when constant $c$ satisfies
$0<c<\frac{2p}{\alpha p+1}-\frac{2}{\alpha+1}=\frac{2(p-1)}{(\alpha+1)(\alpha p+1)}.$
In our current case, we instead obtain the condition
$0<c<\frac{2p}{\alpha p+1}\exp\left(\frac{1}{\alpha}\right)\sqrt{\frac{1}{\alpha}+1}-\frac{2}{\alpha+1},$
which yields a slightly higher upper bound.
Appendix B Experiments
We will, firstly, provide derivations of Lasso and Ridge shrinkage estimators for mean and variance.
Lasso mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda|\mu|\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{1}$ – penalty to the mean and the solution is given by the soft-thresholding operator:
$\hat{\mu}=sign(\bar{x})\max{(0,|\bar{x}|-\frac{\lambda}{2n}),}$
where $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ is the sample mean. This form encourages shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large.
Lasso variance shrinkage. A similar penalized formulation can be used for estimating the variance, leading to a soft-thresholding-type estimator for sample variance $s^{2}$:
$\hat{\sigma}^{2}=\max{(0,s^{2}-\frac{\lambda}{2}),}$
where $\lambda>0$ is regularization hyperparameter, $s^{2}=\frac{1}{n}\sum_{i}(x_{i}-\bar{x})^{2}.$
Ridge mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda\mu^{2}\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{2}$ – penalty to the mean. Taking the derivative and setting it to zero yields the following solution:
$\hat{\mu}=\frac{\sum_{i}x_{i}}{n+\lambda}.$
Ridge variance shrinkage. Similarly, we can obtain the ridge estimate for the variance:
$\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$
where $s^{2}$ is the sample variance and $\lambda>0$ is the regularization parameter.
Now let us discuss reproducibility of experiments on each dataset.
B.1 CIFAR10 dataset
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 4.
We used the standard ResNet-9 architecture, modifying only the batch normalization layers according to the method under consideration. For data splits, we used the standard predefined training and test sets. Training time was approximately $10$ minutes per run, with negligible variation across methods.
All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each experiment was repeated $11$ times to estimate the standard deviation of the performance metrics, and we tested three different batch sizes $(32,64,128)$, resulting in a total training time of approximately $1800$ minutes.
We used SGD with Nesterov acceleration and default parameters, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. We employed early stopping based on validation accuracy on data without noise to prevent overfitting. We used cross-entropy loss for training, and report classification accuracy as the main evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, we selected the regularization parameter $\lambda\in(10^{-3},10^{-1})$ yielding the best performance on validation dataset without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
Standard BN (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
Standard BN (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
Our BN (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our BN (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our BN (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Mean corrected BN (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Mean corrected BN (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
Mean corrected BN (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Khoshsirat et al. BN [15] (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat et al. BN [15] (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat et al. BN [15] (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
Lasso corrected BN (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
Lasso corrected BN (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
Lasso corrected BN (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
Ridge corrected BN(32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
Ridge corrected BN(64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
Ridge corrected BN(128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
Table 4: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
B.2 Cityscapes dataset
We performed segmentaion experiments on Cityscapes dataset – high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. In all our experiments we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 5. We use the HRNetV2-W16 architecture (where 16 denotes the width of the high-resolution convolution) as our baseline model. The original training set was randomly split into $80\%$ for training and $20\%$ for validation, while the original validation set was used as the test set. All images were resized from $1024\times 2048$ to $512\times 1024$ for computational efficiency. To account for variability and estimate the standard deviation of our results, each experiment was repeated $7$ times for each batch size $(16,32).$ Each run took approximately $1.5$ hours on a single NVIDIA A100 GPU with $64$ GB of RAM and $8$ CPU cores, resulting in a total training time of around 126 hours.
We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the mean Intersection over Union (mIoU) score on validation data without noise was employed to prevent overfitting. We used cross-entropy loss for training and report mIoU as the primary evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
Standard BN (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
Our BN (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our BN (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Mean corrected BN (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
Mean corrected BN (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
Khoshsirat et al. BN [15] (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsirat et al. BN [15] (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
Lasso corrected BN (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
Lasso corrected BN (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
Ridge corrected BN(8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
Ridge corrected BN(16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
Table 5: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
B.3 Details of Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN with Stein shrinkage on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). This data was obtained from the PPMI (Parkinson’s Progressive Markers Initiative) portal after a formal request was filled out. Please see data access citation and acknowledgment below for this data.
DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline involving Eddy currents and motion correction etc., using the FSL software [13]. DTIs are matrix-valued images in which each voxel takes on a value that is a $(3,3)$ symmetric positive definite (SPD) matrix representing the covariance matrix of a zero-mean local Gaussian at the voxel. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization is applied to these scalar feature maps to remove site effects caused by distinct acquisition protocols and equipment used. Harmonization was achieved using the method called ComBat adapted to diffusion tensor images and associated software presented in [12]. It should be noted that there are other sophisticated hamronization methods that one can apply to the data but this one is simple. The data set is imbalanced and consists of $490$ (130 PD and 360 Controls) FA maps. To address this, we use mixup-based[31] augmentation with sampling bias favoring the minority class (PD), this approach allowed us to generate additional training examples. We follow the same procedure in our other experiments for the setup and use a 5-fold cross validation to compare 3D CNN models with different variants of 3D BatchNorm layer. The architecture consists of four 3D convolutional layers, each followed by a 3D BatchNorm layer, and three linear layers for binary classification. The 3D input tensors have spatial dimensions of $(128\times 128)$ and a depth of $64$. We conduct experiments using two batch sizes ($10$ and $16$) and optimize the binary cross-entropy loss. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 6. We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the accuracy score on validation data without noise was employed to prevent overfitting. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise. All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each configuration take around $20$ minutes, which end up to the 20 hours for the whole experiments.
Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard BN (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our BN (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our BN (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Mean corrected BN (10) 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
Mean corrected BN (16) 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
Khoshsirat et al. BN [15] (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsirat et al. BN [15] (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso corrected BN (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso corrected BN (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge corrected BN(10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge corrected BN(16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
Table 6: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data. Values depict mean of accuracies $\pm$ standard deviation.

### v2

@@
 Abstract
-Batch normalization (BN) is a ubiquitous operation in deep neural networks used primarily to achieve stability and regularization during network training.
-BN involves feature map centering and scaling using sample means and variances, respectively.
-Since these statistics are being estimated across the feature maps within a batch, this problem is ideally suited for the application of Stein’s shrinkage estimation, which leads to a better, in the mean-squared-error sense, estimate of the mean and variance of the batch.
-In this paper, we prove that the Stein shrinkage estimator for the mean and variance dominates over the sample mean and variance estimators in the presence of adversarial attacks when modeling these attacks using sub-Gaussian distributions.
-This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and use it in image classification (segmentation) tasks with and without adversarial attacks.
-We present SOTA performance results using this Stein corrected batch norm in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
+Batch normalization (BN) is a ubiquitous operation in deep neural networks, primarily used to improve stability and regularization during training.
+BN centers and scales feature maps using sample means and variances, which are naturally suited for Stein’s shrinkage estimation.
+Applying such shrinkage yields more accurate mean and variance estimates of the batch in the mean-squared-error sense.
+In this paper, we prove that the Stein shrinkage estimator of the mean and variance dominates over the sample mean and variance estimators, respectively, in the presence of adversarial attacks modeled using sub-Gaussian distributions.
+Furthermore, by construction, the James–Stein (JS) BN yields a smaller local Lipschitz constant compared to the vanilla BN, implying better regularity properties and potentially improved robustness.
+This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and the use of it in image classification and segmentation tasks with and without adversarial attacks.
+We present SOTA performance results using this Stein-corrected BN in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data, and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
+Keywords: James-Stein shrinkage estimator, Admissibility, Batch normalization, Sub-Gaussian random variables, Adversarial attacks
 1 Introduction
-Deep Neural Networks are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and sciences.
-In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of deep neural networks.
-Deep network training is well known to be time intensive and unstable.
-Batch normalization (BN) is a particularly useful tool in achieving numerical stability and regularization during network training.
-BN typically involves centering and scaling the feature maps.
-The centering and scaling operations involve computing the sample means and variances, respectively, from the samples/members in the batch.
+Deep Neural Networks (DNNs) are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and science.
+In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of DNNs.
+DNN training is well known to be time intensive and unstable.
+Batch normalization (BN) is a particularly useful tool in achieving regularization and numerical stability during network training.
+BN typically involves centering and scaling feature maps.
+The centering and scaling operations involve computing the sample means (SMs) and variances (SVs) from the samples/members in the batch, respectively.
 This task is ideally suited for the application of Stein’s shrinkage method, which is well known in the field of statistics.
-Before delving into the Stein’s shrinkage method, let us briefly review the widely popular batch normalization in deep learning.
-Batch normalization was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift Ioffe-ICML15 .
-Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) ba2016layernormalization and Transformers, spectral normalization in generative adversarial networks (GANs) Miyato-ICLR18 , and many others.
-Improvements over BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training has also been proposed in Zhang_2021_CVPR and batch normalization preconditioning in Lange-JMLR2022 , wherein the pre-conditioner is constructed using batch statistics.
-These normalization algorithms do not address the effectiveness of improvements in stability or convergence of training in the presence of adversarial attacks.
-In this paper, we address the problem of developing a BN that leads to better prediction accuracy from the deep network used in classification in the presence of adversarial attacks.
-Adversarial attacks have been modeled via additive bounded perturbations MauryaArxiv22 .
-Some of the popular adversarial attack methods that are popular and satisfy the bounded perturbation condition are FGSM (fast gradient sign method) szegedy2014 and an iterative variaint of PGD (projected gradient descent).
-Numerous articles exist in the literature that depict ways to overcome a drop in the accuracy of classifiers in the presence of adversarial attacks.
-Our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, however, to adapt an existing shrinkage estimator specifically, the well known James-Stein shrinkage estimator and prove that it retains its dominance over the sample mean and variance in the presence of adversarial attacks.
-Dominance here refers to smaller mean squared error (MSE) – between the estimate and the true value – when compared to the sample mean and variance.
-This smaller MSE implies a better BN in each layer and as we demonstrate empirically, a better stability and prediction accuracy from the deep network.
-By making BN stable in the presence of adversarial attacks this work could have potential applications that might have societal impact.
-1.1 A Brief on the James-Stein Shrinkage Estimator
-In the following we present a brief introduction to the James-Stein (JS) shrinkage estimator and motivate it’s aptness in the context of batch normalization in deep neural networks.
+Before delving into Stein’s shrinkage method, let us briefly review the widely popular BN in DNNs.
+BN was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift (Ioffe and Szegedy, 2015).
+Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) (Jimmy Lei et al., 2016) and Transformers, spectral normalization in generative adversarial networks (GANs) (Miyato et al., 2018), and many others.
+Improvements in BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training have also been proposed in Zhang et al.
+(2021) and BN preconditioning in Susanna et al.
+(2022), wherein the preconditioner is constructed using batch statistics.
+These normalization algorithms do not address the effectiveness of theoretical improvements in the stability or convergence of training in the presence of adversarial attacks.
+The primary motivation of this paper is to explore whether classical statistical techniques, specifically James–Stein (JS) shrinkage estimators, can bring theoretical and practical benefits when applied to DNNs.
+Although many advances in adversarial robustness focus on architectural changes or alternative training objectives, our approach is fundamentally different: we revisit a classical estimator from multivariate statistics and investigate how its well-understood risk-reduction properties behave in the presence of adversarial perturbations.
+This motivation comes from previous empirical work by Khoshsirat and Kambhamettu (2024), which demonstrated that applying shrinkage estimators to the mean and variance of BN layers improved performance under standard non-adversarial conditions.
+The improvement can be attributed to the fact that shrinkage estimators reduce the mean squared error (MSE) under squared loss when estimating multiple parameters jointly – a setting that naturally arises in BN, where each layer must estimate a vector of means and variances simultaneously.
+In this work, we extend this line of inquiry to adversarial scenarios.
+Specifically, we address the problem of designing a variant of BN that maintains or improves the prediction accuracy in the presence of adversarial attacks, modeled as additive bounded perturbations (Maurya and Honorio, 2022).
+Popular attack methods satisfying this bounded perturbation assumption include the fast gradient sign method (FGSM) (Goodfellow et al., 2015) and the iterative variant projected gradient descent (PGD).
+Unlike many existing defenses that focus on robust losses or architectural modifications, our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, but rather, to adapt an existing shrinkage estimator, specifically, the well known JS-shrinkage estimator and prove that it retains its dominance over the SM and SV in the presence of adversarial perturbations.
+We theoretically show and empirically confirm that this dominance translates into more stable BN statistics, which in turn are observed to yield improved robustness and prediction accuracy.
+1.1 A Brief Note on the JS-shrinkage Estimator
+In the following we present a brief introduction to the JS-shrinkage estimator and motivate its aptness in the context of BN in DNNs.
 This will provide the necessary background and motivation for our work presented in this paper.
-Shrinkage estimation is a parameter estimation technique that is primarily used for variance reduction and to improve parameter estimation accuracy by shrinking the parameter towards a fixed target.
-For instance, shrinking the estimate of the mean of a Gaussian toward zero.
-It is especially useful in high-dimensional and scarce data settings.
-The shrinkage, of course, leads to a biased estimate, but with smaller variance, which is sought after in many applications.
-The James-Stein (JS) shrinkage estimator was proposed in the field of statistics as an estimator of the mean for samples drawn from a Gaussian distribution and has been shown to dominate the maximum likelihood estimator (MLE) in terms of the risk (mean-squared error) James-Stein61 .
-Dominance here refers to the risk associated with the JS-estimator being at most as much as that of the MLE and that there exists at least one set of parameter values for which it is strictly less.
-More precisely, if $\mathbf{X}$ is a $p$=dimensional random vector whose mean is to be estimated and it is assumed that $\mathbf{X}\sim N(\mathbf{\mu},\mathbf{I})$, then, $E(||\mathbf{\mu}^{JS}-\mathbf{\mu}||^{2})\leq E(||\mathbf{\mu}^{MLE}-\mathbf{\mu}||^{2})$.
-It is important to note that the JS-estimator improves the overall risk of estimating the mean (parameter) vector and not the risk of the individual elements of the mean vector.
-At the foundation of the JS-shrinkage estimator lies what is known as Stein’s (identity) lemma, which characterizes the normal distribution.
-Stein’s (identity) lemma is also at the foundation of the Stein’s method which is used for distributional comparisons.
-We will not discuss this latter well-known method in statistics here and instead refer the interested reader to a recent comprehensive survey Anastasiou23 .
-Since its inception, the JS-shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family Brown66 , exponential family Hudson78AOS , Gamma distribution DasGupta86AOS , distributions with quadratic varaince function XianchaoXie-Brown-AOS16 and FathiAOS2022 for relaxing the Gaussian assumption in high dimensions.
-For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to Kubokawa24 .
-In addition, more recently, Stein’s shrinkage estimate has been generalized to distributions on manifold-valued random variables Yang-Doss-Vemuri22 .
+[…]
+In the case of the JS-estimator, the parameter $\theta$ is a $p$-dimensional mean vector $\boldsymbol{\mu}$ of a Gaussian random vector $\mathbf{X}\sim N(\boldsymbol{\mu},\mathbf{I})$.
+Under squared error loss, the JS estimator $\boldsymbol{\mu}^{JS}$ dominates the maximum likelihood estimator (MLE) $\boldsymbol{\mu}^{MLE}$, which means that
+$\mathbb{E}\!\left[\|\boldsymbol{\mu}^{JS}-\boldsymbol{\mu}\|^{2}\right]\ \leq\ \mathbb{E}\!\left[\|\boldsymbol{\mu}^{MLE}-\boldsymbol{\mu}\|^{2}\right],$
+with strict inequality for some $\boldsymbol{\mu}$. Importantly, this improvement concerns the overall estimation risk of the mean vector, not the risk of individual components.
+At the foundation of the JS-shrinkage estimator lies Stein’s lemma (also known as Stein’s identity), which characterizes the normal distribution.
+This lemma is also fundamental to Stein’s method, widely used for distributional comparisons.
+We will not discuss this well-known method here and instead refer the interested reader to the recent comprehensive survey by Andreas et al.
+(2023).
+Since its inception, the JS shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family (Lawrence, 1966), exponential family (Harold M., 1978), gamma distribution (Anirban, 1986), distributions with quadratic variance function (Xianchao and Lawrence, 2016) and work by Fathi et al.
+(2022) to relax the Gaussian assumption in high dimensions.
+For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to (Tatsuya, 2024).
+In addition, more recently, Stein’s shrinkage estimate has been generalized to the multiple mean estimation problem in the context of kernel mean embeddings using convex combinations of empirical mean estimators (Blanchard et al., 2024), and to distributions on manifold-valued random variables (Chun-Hao et al., 2024).
 1.2 Relevance of JS-estimator to BN
-Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to batch normalization in deep neural networks and how it can be applied to improve the stability and accuracy of predictions.
-A wide range of normalization techniques have been developed to facilitate the training of deep neural networks by normalizing intermediate feature distributions and improving the stochastic optimization dynamics.
-One of the earliest and most influential approaches is BN, introduced in Ioffe-ICML15 , which normalizes features using statistics computed over the entire mini-batch.
-This method was originally motivated by the desire to reduce internal covariate shift, though subsequent work has highlighted its strong impact on the optimization landscape conf/nips/SanturkarTIM18 as a key factor behind its effectiveness.
-In the following, we give a brief and precise description of different types of normalization that are in vogue in literature.
-Let $X\in\mathbb{R}^{N\times C\times H\times W}$ denote the feature tensor at a given layer of a deep neural network, where $N$ – batch size, $C$ – number of channels (feature maps), $H,W$ – spatial dimension.
-Formally, many normalization techniques follow a common template.
-Each element $x_{i}\in X$ is normalized as: $\hat{x_{i}}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}}$, where $i=(i_{N},i_{C},i_{H},i_{W})$ is a multi-index ranging over the tensor dimensions and $\mu_{i},\sigma^{2}_{i}$ are the empirical mean and variance computed over a subset $S_{i}\subseteq X:$ $\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\ \ \sigma_{i}^{2}=\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2}.$ The choice of $S_{i}$ determines the normalization scheme and the available choices are the following: (i) Batch Normalization (BN)  Ioffe-ICML15 : $S_{i}^{BN}={(n^{\prime},h^{\prime},w^{\prime})|n^{\prime}\in[1,N],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed over all samples in the batch and all spatial positions, but separately for each channel $i_{c}.$ (ii) Layer Normalization (LN)  ba2016layernormalization : $S_{i}^{LN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in[1,C],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed independently for each sample (fixed $i_{N}$) across all channels and spatial locations.
-(iii) Instance Normalization (LN)  ulyanov2017instancenormalizationmissingingredient : $S_{i}^{IN}={(h^{\prime},w^{\prime})|h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Normalization is per-sample and per-channel — only over the spatial dimensions.
-(iv) Group Normalization (GN)  wu2018groupnormalization : $S_{i}^{GN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in C_{\text{group}},h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Channels are split into groups; statistics are computed over spatial dimensions and within the channel group.
-After normalization, a learnable affine transformation is applied to preserve representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
-While normalization methods like LayerNorm, InstanceNorm, and GroupNorm have been introduced to mitigate batch-size dependence, BatchNorm remains a central component in many modern architectures and continues to attract interest both in theory and practice.
-A key observation is that BatchNorm estimates a large number of mean and variance values simultaneously — one pair for each channel.
-This can be naturally viewed as a multitask estimation problem: we are estimating a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$ each based on a finite number of observations.
-In such settings, James–Stein shrinkage estimators offer a statistically grounded way to improve estimation by borrowing strength across tasks — in this case, across channels.
+Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to BN in DNNs and how it can be applied to improve the stability and accuracy of predictions.
+A wide range of normalization techniques have been developed to facilitate the training of DNNs by normalizing intermediate feature distributions and improving the stochastic optimization dynamics.
+One of the earliest and most influential approaches is BN, introduced in Ioffe and Szegedy (2015), which normalizes intermediate features using statistics computed over a mini-batch, estimating for each channel both a mean and a variance.
+In the following, we give a brief and precise description of different types of normalization that are in vogue in the literature.
+Let $X\in\mathbb{R}^{N\times C\times H\times W}$ be the feature tensor, where $N$ is the batch size, $C$ the number of channels (feature maps), and $H,W$ the spatial dimensions.
+Many normalization schemes follow the same basic operation:
+$\hat{x}_{i}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}},\quad\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\quad\sigma_{i}^{2}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2},$
+where $S_{i}$ is the subset of tensor elements over which the statistics are computed.
+The choice of $S_{i}$ defines the normalization type: BN averages over the batch and spatial positions (per channel) (Ioffe and Szegedy, 2015), LayerNorm averages over channels and spatial positions (per sample) (Jimmy Lei et al., 2016), InstanceNorm averages over spatial positions (per sample, per channel) (Ulyanov et al., 2016), and GroupNorm averages over spatial positions and a group of channels (Wu and He, 2018).
+After normalization, a learnable affine transformation is applied to preserve the representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
+Despite the development of alternatives less sensitive to batch size, BN remains widely used due to its strong empirical impact on optimization dynamics (Santurkar et al., 2018) and generalization.
+A key observation is that BN estimates a large number of mean and variance values simultaneously, one pair for each channel.
+This can be naturally viewed as a multitask estimation problem: we estimate a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$, each based on a finite number of observations.
+In such settings, JS-shrinkage estimators offer a statistically grounded way to improve estimation by sharing information across tasks, in this case, across channels (feature maps).
 The classical result suggests that when estimating multiple parameters jointly under squared loss, shrinking individual estimates toward their grand mean can reduce overall error.
-Applying this idea in the context of BatchNorm could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
-Having established adequate motivation for application of Stein’s shrinkage to BN in a deep network, we are now ready to narrate the rest of the story.
+Applying this idea in the context of BN could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
+Having established adequate motivation for application of Stein’s shrinkage to BN in a DNN, we are now ready to narrate the rest of the story.
 But before that, a word on what is to follow.
@@
 In Section 4, we present experimental results that compare the performance of the proposed shrinkage to SOTA on several benchmark data sets.
-Finally in Section 5, we present the discussion and conclusions.
-2 Stein’s Shrinkage Applied to Batch Normalization
-In this section, we apply the Stein’s shrinkage to the BN and present the formulas for calculation of the batch normalization in a deep neural network.
-We will limit ourselves primarily to convolutional neural networks (CNNs) but the formulas can be applied to any architecture with some minor changes.
+Finally, in Section 5, we present the discussion and conclusions.
+2 Stein’s Shrinkage Applied to BN
+In this section, we apply Stein’s shrinkage to the BN and present formulas for it’s computation in DNNs.
 Let $X\sim\mathcal{N}_{p}(\theta,I_{p})$, where $\theta\in\mathbb{R}^{p}$ is an unknown parameter vector and $I_{p}$ is the $p$-dimensional identity matrix.
-The James Stein shrinkage estimator of the unknown parameter $\theta$ is given by, $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X$.
-It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator but is obviously a biased estimator.
+JS-shrinkage estimator of the unknown parameter $\theta$ is given by
+$\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X.$
+It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator (MLE) but is obviously a biased estimator.
 However, a smaller mean squared error is desirable in many applications.
-Most recently, JS-estimator for the calculation of the mean in BN was used in khoshsirat2023improvingnormalizationjamessteinestimator .
-However, the authors in khoshsirat2023improvingnormalizationjamessteinestimator used the same JS-estimator formula to shrink the mean as well as the variance in BN.
-Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the sample variance is not normally distributed but instead follows a scaled chi-squared distribution - a special case of the Gamma distribution.
+Most recently, JS-estimator for the computation of the mean in BN was used in Khoshsirat and Kambhamettu (2024).
+However, the authors in Khoshsirat and Kambhamettu (2024) used the same JS-estimator formula to shrink the mean as well as the variance in BN.
+Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the SV is not normally distributed but instead follows a scaled chi-squared distribution, a special case of the Gamma distribution.
 To address this, we adopt a principled alternative.
-More specifically, for mean correction, we adopt the James-Stein shrinkage under the assumption that individual elements of the feature tensor in the BN are normally distributed and homoscedastic.
-The homoscedasicity assumption simplifies our calculations and we will address relaxing this assumption in future.
-Under this assumption, the sample mean is normally distributed, and the sample variance follows a chi-squared distribution (i.e., a Gamma distribution): $\hat{\sigma}_{i}^{2}\sim\frac{\sigma_{i}^{2}}{n}\cdot\chi^{2}_{n-1}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$, which enables us to apply shrinkage to the scale parameter of the Gamma distribution DasGupta86AOS .
-Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ independently for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$.
-Then JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
-$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right]$ (1)
-To incorporate the James-Stein correction into Batch Normalization, we compute the statistics for each channel and apply the shrinkage formulas as outlined below.
-Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions given by:
+Much of the theoretical work on BN implicitly relies on the assumption that activation maps that form the input to normalization layers are approximately Gaussian (Lee et al., 2018; Neal, 1994b, a).
+This Gaussianity assumption enables analytical tractability and underpins many prior analyses of BN dynamics.
+Following this, we assume that the individual elements of the feature tensor in BN are approximately normally distributed.
+Under this assumption, the SM is normally distributed, while the SV follows a chi-squared distribution.
+Specifically, for a given channel with $n=N\cdot H\cdot W$ observations, $\hat{\sigma}_{i}^{2}\sim(\frac{\sigma_{i}^{2}}{n}\chi^{2}_{n-1})\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$.
+Here, $\chi^{2}$ denotes the $\chi$-squared distribution, which is a special case of the well known Gamma distribution with shape parameter $\alpha$ and scale parameter $\beta$.
+This enables us to apply the shrinkage to the scale parameter of the Gamma distribution using the Stein shrinkage estimator presented in Anirban (1986).
+Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ be independently distributed for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$.
+Then, JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
+$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\quad\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right].$ (1)
+To incorporate the JS correction into BN, we compute the statistics for each channel and apply the shrinkage formulas as outlined below.
+Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions:
 $\mu_{C}=\frac{1}{NHW}\sum_{n,h,w}x_{n,c,h,w},\quad\sigma^{2}_{C}=\frac{1}{NHW}\sum_{n,h,w}(x_{n,c,h,w}-\mu_{C})^{2}.$
-We then apply shrinkage to these estimates. The James-Stein corrected mean and variance are:
+We then apply shrinkage to these estimates. The JS-corrected mean and variance are:
 Stein Shrinkage Forumals for Mean & Variance $\mu^{\mathrm{JS}}_{C}$ $=\left(1-\frac{(c-2)\sigma^{2}_{\mu_{C}}}{\|\mu_{C}\|^{2}_{2}}\right)\mu_{C},\quad{\sigma_{C}^{2}}^{\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+\tilde{c}\cdot V,$ (2) $\text{where }V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad\tilde{c}\in\left[0,\dfrac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$
@@
 Final Normalized (scaled and shifted) Output $\hat{x}=\frac{x-\mu^{\mathrm{JS}}}{\sqrt{{\sigma^{2}}^{\mathrm{JS}}+\epsilon}},\quad y=\gamma\hat{x}+\beta.$ (3)
-Equations 2 and 3 respectively are the formulas we advocate for use in BN corrections and will be used in all our experiments.
-We are now ready to present the theoretical results of dominance of these shrinkage estimators in the presence of adverserial attacks.
-3 Dominance of Stein Shrinkage in the Presence of Adverserial Attacks
-In this section, we present the main theoretical contribution of our work: the James–Stein estimator continues to exhibit dominance even under adversarial perturbations.
+Equations 2 and 3, respectively, are the formulas we advocate for use in BN corrections and will be used in all our experiments.
+We would like to mention that the choice of the shrinkage constant can, in principle, influence the estimator.
+When the constant approaches zero, the estimator reduces to the empirical variance, so values near the boundary of the admissible interval are of primary theoretical interest.
+In our experiments, we did not observe a significant sensitivity to moderate changes in this constant.
+Following the precedent set in the original work on shrinkage estimators for the gamma distribution (Anirban, 1986), we adopted boundary values of the shrinkage constant in our experiments, as detailed in Section 4.
+We are now ready to present the theoretical results of the dominance of these shrinkage estimators in the presence of adversarial attacks.
+3 Dominance of Stein Shrinkage in the Presence of Adversarial Attacks
+In this section, we present the main theoretical contribution of our work namely, that the JS-estimator continues to exhibit its dominance even under adversarial perturbations.
 We begin by formalizing how adversarial attacks can be modeled probabilistically.
-Then, we introduce the necessary preliminaries and outline the main ideas behind the proof.
-We model adversarial perturbations using a sub-Gaussian distribution, which is well-motivated both practically and theoretically.
-Adversarial perturbations are typically bounded in magnitude (e.g., under an $l_{p}$-norm constraint) to ensure the model remains learnable — a standard assumption in the literature cohen2019certifiedadversarialrobustnessrandomized ; zhai2019adversariallyrobustgeneralizationjust .
+Then, we introduce the necessary preliminaries and outline the main ideas behind the proof and include the detailed proof in the Appendix A.
+We model adversarial perturbations using a sub-Gaussian distribution, which is well motivated both practically and theoretically.
+Adversarial perturbations are typically bounded in magnitude (for example, under an $l_{p}$-norm constraint) to ensure that the model remains learnable - a standard assumption in the literature (Cohen et al., 2019; Schmidt et al., 2018).
 Since bounded random variables are inherently sub-Gaussian, this provides a natural modeling choice.
 Additionally, the sub-Gaussian model enables a tractable analysis via the use of concentration inequalities and allows one to work with a well-defined variance proxy.
-Following prior workMauryaArxiv22 , we assume that each adversarial perturbation vector has coordinates bounded in $[-\varepsilon,\varepsilon]$ and can be represented as a zero-mean sub-Gaussian vector with variance proxy $2\varepsilon^{2}$: $Y\sim\text{subG}(2\varepsilon^{2}),\quad\text{i.e., for all }v\in\mathbb{R}^{p}\text{ with }\|v\|=1,\ \ \text{and}\ \ \mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\frac{2\varepsilon^{2}\lambda^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$ This assumption allows us to model the observed vector as the sum of noise-free data and adversarial noise: $Z=X+Y,\ \ \ X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$ For simplicity we assume that $X$ and $Y$ are independent, which is currently a limitation of our model.
-Other noise models that are data dependent will be considered in our future work.
+Following (Maurya and Honorio, 2022), we assume that adversarial perturbations $Y\in\mathbb{R}^{p}$ are mean-zero sub-Gaussian with variance proxy $2\varepsilon^{2}$, i.e., $\mathbb{E}\!\left[e^{\lambda\langle v,Y\rangle}\right]\leq e^{\varepsilon^{2}\lambda^{2}},\ \forall\lambda\in\mathbb{R},\ \forall v\in\mathbb{R}^{p}:\|v\|=1,$ which we denote compactly as $Y\sim\mathrm{SG}(2\varepsilon^{2})$.
+This assumption allows us to model the observed vector as a sum of noiseless data and adversarial noise: $Z=X+Y,\quad X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$
 3.1 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Normally Distributed Random Variables
-Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\in\mathbb{R}^{p}$ is independent, mean-zero sub-Gaussian vector with a known proxy variance $2\varepsilon^{2}.$ Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{\mathrm{MLE}}$ and the JS-estimator
+Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\sim\mathrm{SG}(2\varepsilon^{2})$.
+Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{0}=Z$ and the JS-estimator
 $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{(p-2)\hat{\sigma}^{2}}{\|Z\|^{2}_{2}}\right)Z.$
 under squared error loss.
-It is well-known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{\mathrm{MLE}}$ when $Y=0$ and $p\geq 3$ James-Stein61 .
+It is well known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{0}$ when $Y=0$ and $p\geq 3$ (James and Stein, 1961).
 We aim to show that this dominance remains valid even under additive sub-Gaussian noise, provided the proxy variance $2\varepsilon^{2}$ is sufficiently small.
@@
 JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gaussian Model
-Theorem 3.1.
+Theorem 1
 Let $\theta\in\mathbb{R}^{p}$ be an unknown parameter vector.
-Suppose we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent, mean-zero sub-Gaussian vector with parameter $2\varepsilon^{2}$, i.e., for all unit vectors $v\in\mathbb{R}^{p}$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\varepsilon^{2}\lambda^{2}\right)$.
-Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the maximum likelihood estimator by $\hat{\theta}^{\mathrm{MLE}}=Z$.
-Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the MLE in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{MLE}},\theta).$
-We now provide a sketch of the proof and refer the reader to the Appendix for more details.
-Sketch of Proof.
-We can rewrite the risk of the JS-estimator as:
-$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
+Suppose that we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent $Y\sim\mathrm{SG}(2\varepsilon^{2})$.
+Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the observation-based estimator by $\hat{\theta}^{0}=Z$.
+Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the observation-based estimator in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{0}},\theta).$
+We now provide a sketch of the proof for this theorem and refer the reader to the Appendix A for a detailed proof.
+Proof [Sketch of the Proof] We can rewrite the risk of the JS-estimator as:
+$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{0}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
 The last two terms in equation 4 can be rewritten as:
 $-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\right)$
-Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that: $\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$ Hence, the risk of the James–Stein estimator is strictly less than that of the MLE for $p\geq 3$ completing the proof.
-∎
+Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that:
+$\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$
+Hence, the risk of the JS-estimator is strictly less than that of the observation-based estimator for $p\geq 3$, thus completing the proof.
 We are now ready to address the issue of shrinkage estimation of the variance parameter in the above normally distributed random variable with additive sub-Gaussian perturbation.
 3.2 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Gamma Distributed Random Variables
-In many applications, empirical variances follow a scaled chi-squared distribution, which is a special case of the gamma distribution.
-Given observations $x_{ij}$, define the empirical variance: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$.
+In many applications, SVs follow a scaled chi-squared distribution, which is a special case of the gamma distribution.
+Given observations $x_{ij}$, define the SV: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$.
 Thus, estimating $\sigma_{i}^{2}$ reduces to estimating the scale parameter $\beta_{i}$ in a Gamma distribution with fixed shape parameter $\alpha$.
-For this task, we can use the following theorem from literature:
-Theorem 3.2 (JS-Estimator Shrinkage Dominance for the Gamma Model DasGupta86AOS ).
+For this task, we can use the following theorem from literature (Anirban, 1986):
+Theorem 2 (Dominance of the JS-Estimator for the Gamma Model)
+Let $X_{1},\dots,X_{p}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ be independently distributed.
+Consider the naive estimator $\hat{\beta}_{i}^{\,0}=\frac{X_{i}}{\alpha+1},i=1,\dots,p.$ Define the JS-type shrinkage estimator $\hat{\beta}_{i}^{\,\mathrm{JS}}=\hat{\beta}_{i}^{\,0}+cV,V=\Biggl(\prod_{j=1}^{p}X_{j}\Biggr)^{\!1/p}.$ If $c\in\left[\,0,\;\frac{2(p-1)}{(\alpha+1)(\alpha p+1)}\,\right],$ then the shrinkage estimator dominates the naive estimator in risk, i.e.
+$R\!\left(\hat{\beta}^{\,\mathrm{JS}},\beta\right)\;\leq\;R\!\left(\hat{\beta}^{\,0},\beta\right),\forall\;\beta\in\mathbb{R}_{+}^{p}.$
+Applying the previous theorem to SVs, we obtain the JS-type shrinkage estimator
+$\hat{\sigma}_{i}^{2,\,\mathrm{JS}}\;=\;\frac{n}{n+1}\,\hat{\sigma}_{i}^{2}\;+\;c\,V,\qquad V\;=\;\Biggl(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\Biggr)^{\!1/p},\qquad c\;\in\;\left[\,0,\;\frac{4n(p-1)}{(n+1)\big((n-1)p+2\big)}\,\right].$
+We now present our second key result, namely, that this dominance persists even under additive sub-Gaussian noise for the Gamma-distributed random variables.
+JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gamma Model
+Theorem 3
 […]
-Now, applying the above theorem to empirical variances, we obtain: $\hat{\sigma}_{i}^{2,\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+cV,\quad V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad c\in\left[0,\frac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$ We now present our second key result namely, that this dominance persists under additive sub-Gaussian noise for the Gamma distributed random variable.
-JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gamma Model
-Theorem 3.3.
-[…]
-Remark 3.4.
+Remark 4
 The admissible range for $c$ under sub-Gaussian perturbations includes the classical interval, so standard shrinkage remains valid even in noisy settings.
-Sketch of Proof.
 […]
 We begin by expressing the risk of $\hat{\beta}^{\,\mathrm{JS}}$ as:
-[…]
-[…]
-Thus, the difference in risks depends on the sign of the second term in the above expression expressed as: $c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$ To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term.
-We consider two main cases:
+$R(\hat{\beta}^{\,\mathrm{JS}},\beta)=\mathbb{E}\left(\left\|\hat{\beta}^{\,\mathrm{JS}}-\beta\right\|_{2}^{2}\right)=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}+cV-\beta\right\|_{2}^{2}\right)=$
+$=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}-\beta\right\|_{2}^{2}\right)+2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)\cdot cV\right)+p\mathbb{E}\left(c^{2}V^{2}\right)$
+Thus, the difference in risks depends on the sign of the second term in the above expression expressed as:
+$c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$
+To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term. We consider two main cases:
   • Case 1:
     $c<0$ and
-    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (5)
+    $2\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>0$ (5)
   • Case 2:
     $c>0$ and
-    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (6)
+    $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<0$ (6)
 In both cases outlined above, the proof fundamentally relies on an application of Stein’s lemma for the Gamma distribution given below.
 Since each $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ with known shape and rate parameters, we can invoke the Stein’s identity given in the following lemma:
-Lemma 3.5 (Stein’s Lemma for the Gamma Distribution).
-Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty$.
-Then the following (Stein’s) identity holds: $\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)]$.
+Lemma 5 (Stein’s Lemma for the Gamma Distribution)
+Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that
+$\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty.$
+Then the following (Stein’s) identity holds:
+$\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)].$
 We apply Stein’s identity to the function $h(x)=V$, where $V$ is the geometric mean defined by $V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}$.
-Using the Stein’s identity for the Gamma distributed random variable, which is given by, $\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$ we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get: $\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
-Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality.
+Using the Stein’s identity for the Gamma distributed random variable, which is given by,
+$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$
+we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
+$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
+Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality respectively.
 These tools allow us to bound the expressions for a given choice of $c$ and derive the dominance for the JS-estimator of the variance parameter of the normally distributed random variable under sub-Gaussian perturbations.
-Further technical details and precise bounds are deferred to the appendix. ∎
+Further technical details and precise bounds are deferred to the appendix.
 4 Experiments
 In this section, we present several experimental results demonstrating the performance of BN with our proposed shrinkage estimators in both image classification and segmentation tasks, under standard and adversarial settings.
-To simulate adversarial attacks, we added noise samples drawn from a sub-Gaussian distribution with a fixed variance.
-The sub-Gaussian density is given by $f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$ derived from a mixture of Lévy and Gaussian distributions, and the samples were drawn using the inverse transform sampling technique.
-GEORGIOU20063061 .
-For classification, we employed the ResNet-9 he2015deepresiduallearningimage architecture on the CIFAR-10  cifar10 dataset, as well as a 3D CNN model on the PPMI dataset ppmi-data .
-For object segmentation, we used the HRNetV2 wang2020deep model applied to the Cityscapes datasetCordts2015Cvprw .
+To simulate adversarial attacks, we generated samples of the noise signal modeled by a sub-Gaussian–like distribution with the density
+$f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$
+which arises from a mixture of Lévy and Gaussian distributions (Panayiotis G. and Chris, 2006). The samples were drawn using inverse transform sampling.
+For classification, we employed the ResNet-9 (He et al., 2016) architecture on the CIFAR-10 dataset (Alex et al., 2009), as well as a 3D CNN model on the PPMI dataset (1).
+For object segmentation, we used the HRNetV2 (Wang et al., 2020) model applied to the Cityscapes dataset (Cordts et al., 2015).
 Since BN is known to become unstable when the batch size is small, we report results for varying batch sizes and sub-Gaussian noise levels.
@@
 We remind the reader that our goal here is not to improve the SOTA results on these data in the literature, but to demonstrate the performance gains from the use of BN in the presence (absence) of adversarial attacks with the correct use of Stein shrinkage compared to vanilla BN and other forms of shrinkage applied to BN.
-Throughout all the experiments, we avoided training-time tricks and focused on obtaining a honest comparison.
-BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
-Standard (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
-Standard (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
-Standard (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
-Our (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
-Our (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
-Our (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
-Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
-Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
-Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
-Lasso (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
-Lasso (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
-Lasso corrected (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
-Ridge (32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
-Ridge (64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
-Ridge (128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
-Table 1: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
-In the experiments, we compare our approach mainly with that of Khoshsirat et al.
-khoshsirat2023improvingnormalizationjamessteinestimator , aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes.
+Throughout all experiments, we avoided training-time tricks and focused on obtaining an honest comparison.
+In the experiments, we compare our approach mainly with Khoshsirat and Kambhamettu (2024), aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes.
 In addition, with the correct variance shrinkage, one can achieve higher accuracy in the presence of adversarial perturbations.
 We also tested a variant using only the mean shrinkage estimator and no shrinkage on variance.
-Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are commonly used shrinkage methods.
-Detailed formulas for these estimators and complete experimental setup are provided in the Appendix B.
+Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are classical shrinkage methods in literature.
+Below we summarize the forms of these shrinkage estimators and provide explicit formulas for convenience.
+Derivations can be found, for example, in Hastie et al.
+(2009).
+Lasso mean and variance shrinkage.
+These estimators are obtained by applying an $\ell_{1}$ penalty in penalized least squares problems (Hastie et al.
+(2009)).
+For the mean, this leads to the soft-thresholding solution $\hat{\mu}=\operatorname{sign}(\bar{x})\max\!\left(0,|\bar{x}|-\tfrac{\lambda}{2n}\right),$ where $\bar{x}=\tfrac{1}{n}\sum_{i=1}^{n}x_{i}$ is the SM.
+This form encourages the shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large.
+For variance, an analogous penalized formulation yields $\hat{\sigma}^{2}=\max\!\left(0,\,s^{2}-\tfrac{\lambda}{2}\right),$ where $\lambda>0$ is the regularization hyperparameter and $s^{2}=\tfrac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$ is the SV.
+Ridge mean and variance shrinkage.
+These estimators are obtained by applying an $\ell_{2}$ penalty in penalized least squares problems (Hoerl and Kennard (1970); Hastie et al.
+(2009)).
+For the mean, this yields $\hat{\mu}=\frac{\sum_{i=1}^{n}x_{i}}{n+\lambda},$ and for the variance $\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$ where $s^{2}=\tfrac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$ is the SV and $\lambda>0$ is the regularization parameter.
+Complete experimental setup are provided in the Appendix B.
 Note that in all the tables depicting the results, we have highlighted the best performing BN method (ours) for small batch sizes that are preferred from a computational efficiency perspective as well as when data are scarce.
+Code available at https://github.com/sivolgina/shrinkage.
 4.1 Application to Cifar10 Data
@@
 The results in Table 1 confirm that our proposed BN significantly outperforms the baselines in noisy settings across all batch sizes.
-Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in khoshsirat2023improvingnormalizationjamessteinestimator , our approach shows consistent improvement due to the correct development and use of variance shrinkage.
-The method with only mean correction (see Appendix) performs worse than our full model, highlighting the importance of variance correction.
+Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in Khoshsirat and Kambhamettu (2024), our approach shows consistent improvement due to the correct development and use of variance shrinkage.
+The method with only mean correction performs worse than our full model, highlighting the importance of variance correction.
 Finally, Lasso and Ridge-based approaches do not improve noise performance, showing behavior similar to the standard BN.
+Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
+Standard (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
+Standard (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
+Standard (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
+Ours (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
+Ours (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
+Ours (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
+Mean corrected (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
+Mean corrected (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
+Mean corrected (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
+Khoshsirat et al. (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
+Khoshsirat et al. (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
+Khoshsirat et al. (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
+Lasso (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
+Lasso (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
+Lasso (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
+Ridge (32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
+Ridge (64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
+Ridge (128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
+Table 1: Performance of ResNet-9 with different BN methods under varying subgaussian noise and batch sizes respectively.
+Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data.
+Numbers inside parentheses in the far left column denote batch size.
 4.2 Application to Cityscapes Data
-BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
+Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
 Standard (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
 Standard (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
-Our (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
-Our (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
-Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
-Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
+Ours (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
+Ours (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
+Mean corrected (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
+Mean corrected (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
+Khoshsirat et al. (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
+Khoshsirat et al. (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
 Lasso (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
@@
 Ridge (16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
-Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
-To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset—a more challenging task involving high-resolution urban scene segmentation with fine-grained object boundaries.
+Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively.
+Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data.
+Numbers inside paranthesis in the far left column denote batch size.
+To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset – a more challenging task that involves high resolution urban scene segmentation with fine-grained object boundaries.
 The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation.
-As in the previous experiment, we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics.
+We use the mean intersection over union (mIoU), a standard metric used for semantic segmentation in the computer vision community.
+As in the previous experiment, we injected sub-gaussian noise into the feature maps during validation and testing to simulate corrupted statistics.
 This setup is particularly relevant for segmentation, where maintaining accurate local feature distributions is crucial for pixel-level predictions.
 As shown in Table 2, standard BN fails to retain meaningful performance even under mild noise (e.g., below $10$ mIoU score at $3\%$ corruption).
-In contrast, our proposed method and its mean corrected variant (see the appendix) significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings.
+In contrast, our proposed method and its mean corrected variant significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings.
 Regularization-based approaches (Lasso and Ridge) show negligible improvement under noise and have a behavior similar to that of standard BN.
@@
 In this experiment, the goal is to demonstrate the performance of our BN method on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls).
-DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged.
-The PPMI data was subjected to a standard pre-processing pipeline (see Appendix B) using the FSL software FSLpack .
-DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local Gaussian fit.
+DTI is a noninvasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged.
+The PPMI data were subjected to a standard preprocessing pipeline (see Appendix B).
+using the FSL software (Jenkinson et al., 2012).
+DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local (voxel-level) Gaussian fit.
 These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps.
 The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI.
-Harmonization FORTIN2017149 is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used.
+Harmonization (Jean-Philippe et al., 2017) is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used.
 The data set is highly imbalanced and contains FA maps from 360 Control and 130 PD brain scans.
-This imbalance is addressed using mixup-basedDBLP:journals/corr/abs-1710-09412 augmentation with sampling bias favoring the smaller sized class (PD).
+This imbalance is addressed using mixup-based augmentation (Zhang et al., 2018) with sampling bias favoring the smaller sized class (PD).
 Using the same setup as in previous experiments and a 5-fold cross-validation to compare 3D CNN with different variants of 3D BN, we get results summarized in the Table 3.
 These results further support our earlier observations, even in this challenging classification task.
-BN Method 3% noise 10% noise 15% noise 20% noise 30% noise
-Standard (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
-Standard (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
-Our (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
-Our (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
-Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
-Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
-Lasso (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
-Lasso (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
-Ridge (10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
-Ridge (16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
-Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on PPMI data. Values depict mean accuracies $\pm$ standard deviation.
-5 Conclusions
-In this paper, we presented a novel theory involving the proof of dominance of the Stein’s shrinkage estimators of mean and variance of a Gaussian random variable in the presence of adversarial attacks.
-We applied Stein’s shirnkage to BN in a variety of deep networks and demonstrated superior performance, in comparison to other BN methods, on three benchmark data sets.
-One possible explanation for the improved performance is the connection between model performance under attack and the Lipschitz properties of the network DBLP:journals/corr/abs-2003-02460 .
-Since BN is known to affect the Lipschitz constant conf/nips/SanturkarTIM18 , it would be interesting to investigate the influence of our BN method on this aspect in future work.
-6 Acknowledgments
-This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099. (Corresponding author: Baba C. Vemuri.)
+Since the image data in this experiment are 3D volumetric images, the computational needs are much higher than when handling 2D image data sets, and we therefore used smaller batch sizes in the reported results.
+This helps reduce the computational footprint of the experiment.
+Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
+Standard (10) 94.9 $\pm$ 1.3 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
+Standard (16) 93.3 $\pm$ 1.5 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
+Ours (10) 95.1 $\pm$ 1.7 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
+Ours (16) 95.1 $\pm$ 1.8 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
+Mean corrected (10) 95.3 $\pm$ 2.6 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
+Mean corrected (16) 93.1 $\pm$ 2.4 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
+Khoshsirat et al. (10) 95.1 $\pm$ 2.4 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
+Khoshsirat et al. (16) 94.7 $\pm$ 2.5 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
+Lasso (10) 96.5 $\pm$ 3.0 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
+Lasso (16) 95.1 $\pm$ 1.9 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
+Ridge (10) 95.7 $\pm$ 1.7 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
+Ridge (16) 97.1 $\pm$ 1.3 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
+Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data.
+Values depict mean of accuracies $\pm$ standard deviation.
+Numbers inside paranthesis in the far left column denote batch size.
+4.4 Experiments using Data Dependent Attacks
+To validate our theoretical findings, we conducted a series of experiments using well-known adversarial attacks.
+While numerous attack methods have been proposed, our focus here is on confirming the theoretical insights rather than providing an extensive empirical comparison.
+Therefore, we restricted our study to two of the most widely used attacks, FGSM (Goodfellow et al., 2015) and it’s iterative version PGD.
+The results are presented below; for detailed experimental setups, we refer the reader to the appendix.
+We first evaluate robustness against the FGSM attack.
+As before, we use a ResNet-9 model trained on the CIFAR-10 dataset and HRNetV2 model on the Cityscapes dataset, with the same hyperparameters as in the previous experiments.
+For consistency, we consider batch sizes matching those used in the earlier experiments and run each configuration five times.
+The results at different perturbation levels ($\epsilon=0.03,0.1,0.15,0.2,0.3$) are reported in Tables 4, 5.
+We omit regularized BN baselines on CIFAR-10, as they perform similarly to standard BN.
+BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
+Standard (32) 0.26 $\pm$ 0.04 0.07 $\pm$ 0.05 0.18 $\pm$ 0.18 0.45 $\pm$ 0.53 1.0 $\pm$ 0.96
+[…]
+Standard (128) 0.05 $\pm$ 0.02 0.06 $\pm$ 0.05 0.05 $\pm$ 0.02 0.11 $\pm$ 0.04 0.39 $\pm$ 0.13
+Ours (32) 42.72 $\pm$ 1.16 26.06 $\pm$ 1.5 21.66 $\pm$ 1.31 18.74 $\pm$ 0.99 14.868 $\pm$ 0.94
+Ours (64) 43.29 $\pm$ 0.56 26.17 $\pm$ 0.41 21.66 $\pm$ 0.34 18.69 $\pm$ 0.39 14.8 $\pm$ 0.35
+Ours (128) 36.36 $\pm$ 1.38 20.27 $\pm$ 1.35 16.82 $\pm$ 1.27 14.41 $\pm$ 1.16 11.45 $\pm$ 0.97
+Mean corrected (32) 34.39 $\pm$ 0.84 20.07 $\pm$ 0.72 16.53 $\pm$ 0.76 14.28 $\pm$ 0.6 11.42 $\pm$ 0.49
+Mean corrected (64) 26.82 $\pm$ 1.08 14.4 $\pm$ 1.16 11.99 $\pm$ 0.99 10.29 $\pm$ 0.84 7.98 $\pm$ 0.59
+Mean corrected (128) 16.57 $\pm$ 0.96 8.43 $\pm$ 1.19 7.03 $\pm$ 1.17 6.14 $\pm$ 1.0 4.92 $\pm$ 0.83
+Khoshsirat et al. (32) 37.04 $\pm$ 0.71 22.78 $\pm$ 0.92 19.06 $\pm$ 0.86 16.35 $\pm$ 0.66 12.75 $\pm$ 0.49
+Khoshsirat et al. (64) 38.27 $\pm$ 1.01 22.97 $\pm$ 0.49 19.07 $\pm$ 0.62 16.24 $\pm$ 0.73 12.3 $\pm$ 0.64
+Khoshsirat et al. (128) 33.45 $\pm$ 0.79 18.49 $\pm$ 0.82 15.22 $\pm$ 0.71 12.94 $\pm$ 0.54 9.9 $\pm$ 0.85
+Table 4: Performance of ResNet-9 with different BN methods under varying FGSM attack and batch sizes respectively.
+Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data.
+Numbers inside the parentheses in the far left column denote batch size.
+BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
+Standard (8) 9.4 $\pm$ 1.5 7.6 $\pm$ 0.9 6.2 $\pm$ 1.7 4.9 $\pm$ 2.3 3.8 $\pm$ 2.6
+Standard (16) 10.2 $\pm$ 1.0 5.5 $\pm$ 1.5 4.7 $\pm$ 1.7 4.3 $\pm$ 1.9 3.9 $\pm$ 2.0
+Ours (8) 38.1 $\pm$ 1.4 28.8 $\pm$ 1.8 25.9 $\pm$ 2.4 23.6 $\pm$ 3.2 20.9 $\pm$ 4.5
+Ours (16) 35.4 $\pm$ 0.6 24.9 $\pm$ 1.0 21.8 $\pm$ 1.3 19.8 $\pm$ 1.6 17.3 $\pm$ 1.8
+Mean corrected (8) 33.7 $\pm$ 1.9 24.4 $\pm$ 1.4 22.3 $\pm$ 1.2 20.8 $\pm$ 1.3 18.9 $\pm$ 1.8
+Mean corrected (16) 33.5 $\pm$ 2.7 24.6 $\pm$ 3.3 21 $\pm$ 3.1 19.4 $\pm$ 2.9 17 $\pm$ 3.3
+Khoshsirat et al. (8) 35.9 $\pm$ 2.2 26.2 $\pm$ 2.6 23.4 $\pm$ 2.6 21.7 $\pm$ 2.6 19.7 $\pm$ 2.4
+Khoshsirat et al. (16) 33.4 $\pm$ 1.8 22.9 $\pm$ 2.3 20.4 $\pm$ 2.4 18.87 $\pm$ 2.3 16.9 $\pm$ 2.1
+Lasso (8) 10.5 $\pm$ 5.7 5.7 $\pm$ 3.4 5.0 $\pm$ 2.9 4.8 $\pm$ 2.7 4.8 $\pm$ 2.8
+Lasso (16) 6.8 $\pm$ 2.9 3.6 $\pm$ 1.8 3.3 $\pm$ 1.7 3.5 $\pm$ 1.7 4.2 $\pm$ 2.3
+Ridge (8) 9.7 $\pm$ 1.8 6.2 $\pm$ 0.8 4.8 $\pm$ 0.8 4.3 $\pm$ 0.7 4.1 $\pm$ 0.8
+Ridge (16) 12.9 $\pm$ 6.3 7.0 $\pm$ 1.0 5.9 $\pm$ 0.9 5.5 $\pm$ 0.9 4.9 $\pm$ 1.0
+Table 5: Performance of HRNetV2 with different BN methods under varying FGSM attack and batch sizes respectively.
+Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data.
+Numbers inside the parentheses in the far left column denote batch size.
+In addition, we evaluated robustness under the PGD attack. The experimental configuration mirrors that of the FGSM case; The corresponding results are summarized in Tables 6, 7.
+BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
+Standard (32) 0.08 $\pm$ 0.02 0.08 $\pm$ 0.01 0.0 0.0 0.0
+Standard (64) 0.19 $\pm$ 0.03 0.0 0.0 0.0 0.0
+Standard (128) 0.0 0.0 0.0 0.0 0.0
+Ours (32) 27.46 $\pm$ 0.74 17.27 $\pm$ 0.48 13.5 $\pm$ 0.81 10.74 $\pm$ 0.64 6.8 $\pm$ 0.47
+Ours (64) 32.05 $\pm$ 1.86 21.7 $\pm$ 1.5 17.01 $\pm$ 1.4 13.31 $\pm$ 1.18 8.44 $\pm$ 0.69
+Ours (128) 24.54 $\pm$ 1.7 14.86 $\pm$ 1.54 11.29 $\pm$ 1.39 8.88 $\pm$ 1.08 5.6 $\pm$ 0.5
+Mean corrected (32) 19.57 $\pm$ 1.61 10.84 $\pm$ 1.73 7.83 $\pm$ 1.67 5.98 $\pm$ 1.38 3.52 $\pm$ 0.94
+Mean corrected (64) 13.65 $\pm$ 1.93 7.03 $\pm$ 1.93 4.96 $\pm$ 1.46 3.7 $\pm$ 1.24 2.01 $\pm$ 0.81
+Mean corrected (128) 5.3 $\pm$ 0.71 2.05 $\pm$ 0.38 1.45 $\pm$ 0.27 0.94 $\pm$ 0.23 0.51 $\pm$ 0.11
+Khoshsirat et al. (32) 28.36 $\pm$ 1.44 20.302 $\pm$ 1.6 14.61 $\pm$ 1.27 10.37 $\pm$ 1.15 5.3 $\pm$ 0.6
+Khoshsirat et al. (64) 28.62 $\pm$ 1.08 20.45 $\pm$ 1.02 14.76 $\pm$ 1.15 10.37 $\pm$ 0.64 5.25 $\pm$ 0.42
+Khoshsirat et al. (128) 25.5 $\pm$ 0.47 17.74 $\pm$ 0.54 12.53 $\pm$ 0.04 7.83 $\pm$ 0.14 3.95 $\pm$ 0.2
+Table 6: Performance of ResNet-9 with different BN methods under varying PGD attack and batch sizes respectively.
+Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data.
+Numbers inside the parentheses in the far left column denote batch size.
+BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
+Standard (8) 1.8 $\pm$ 1.4 0.8 $\pm$ 0.7 0.6 $\pm$ 0.5 0.5 $\pm$ 0.3 0.6 $\pm$ 0.4
+Standard (16) 2.0 $\pm$ 1.0 0.5 $\pm$ 0.2 0.3 $\pm$ 0.1 0.2 $\pm$ 0.1 0.2 $\pm$ 0.0
+Ours (8) 10.4 $\pm$ 1.1 4.7 $\pm$ 1.1 3.4 $\pm$ 1.1 2.7 $\pm$ 1.1 2.2 $\pm$ 1.1
+Ours (16) 10.6 $\pm$ 0.9 5.0 $\pm$ 0.8 3.7 $\pm$ 0.8 2.9 $\pm$ 0.8 2.2 $\pm$ 0.6
+Mean corrected (8) 10.1 $\pm$ 0.9 3.7 $\pm$ 0.6 2.4 $\pm$ 0.5 1.8 $\pm$ 0.4 1.3 $\pm$ 0.4
+Mean corrected (16) 10.5 $\pm$ 1.9 4.8 $\pm$ 1.3 3.4 $\pm$ 1.1 2.7 $\pm$ 0.9 2.0 $\pm$ 0.7
+Khoshsirat et al. (8) 8.5 $\pm$ 1.0 2.6 $\pm$ 0.2 1.7 $\pm$ 0.2 1.4 $\pm$ 0.2 1.1 $\pm$ 0.2
+Khoshsirat et al. (16) 9.1 $\pm$ 1.6 3.3 $\pm$ 1.0 2.3 $\pm$ 0.6 1.8 $\pm$ 0.4 1.4 $\pm$ 0.2
+Lasso (8) 2.8 $\pm$ 1.4 2.0 $\pm$ 1.8 1.9 $\pm$ 2.0 1.8 $\pm$ 2.2 1.6 $\pm$ 2.3
+Lasso (16) 2.6 $\pm$ 1.5 1.6 $\pm$ 1.0 1.4 $\pm$ 0.9 1.3 $\pm$ 1.0 1.1 $\pm$ 1.0
+Ridge (8) 2.6 $\pm$ 1.8 1.4 $\pm$ 1.8 1.1 $\pm$ 1.6 0.9 $\pm$ 1.2 0.5 $\pm$ 0.5
+Ridge (16) 2.8 $\pm$ 0.6 1.0 $\pm$ 0.4 0.9 $\pm$ 0.6 0.8 $\pm$ 0.7 0.7 $\pm$ 0.8
+Table 7: Performance of HRNetV2 with different BN methods under varying PGD attack and batch sizes respectively.
+Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data.
+Numbers inside the parentheses in the far left column denote batch size.
+Overall, the findings support our theoretical results: Applying JS-based shrinkage to both the mean and variance yields significantly improved robustness under adversarial settings compared to other BN methods.
+5 Discussion and Conclusion
+One possible explanation for the improved performance of JS-BN under adversarial perturbations lies in its effect on the Lipschitz properties of the network.
+It is well known that smaller local Lipschitz constants are associated with more regular functions, which are inherently more stable to perturbations (Evans and Gariepy, 2015).
+From the empirical side, recent work has shown that local Lipschitzness enforces both robustness and accuracy (Yang et al., 2020).
+Our theoretical results show that JS-BN reduces the local Lipschitz constant relative to standard BN.
+Intuitively, this shrinkage effect smooths the geometry of the loss landscape.
+For BN, we have the following.
+$f(x)=\gamma\frac{x-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta,$
+which yields
+$|f(x)-f(y)|=\left|\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}(x-y)\right|.$
+Thus, the corresponding Lipschitz constant is
+$L=\frac{|\gamma|}{\sqrt{\sigma^{2}+\epsilon}}.$
+Since, by construction, the JS-shrinkage produces a larger variance estimate, $\sigma^{2,\,\mathrm{JS}}>\sigma^{2}$, the denominator increases, resulting in a smaller local Lipschitz constant.
+This naturally leads to an improved smoothness and robustness of the model.
+In summary, in this paper, we present a novel theory that demonstrates the dominance of Stein’s shrinkage estimators for mean and variance over SM and SV under adversarial attacks.
+We applied JS-BN in a variety of DNNs and demonstrated superior performance compared to other BN methods on three benchmark datasets.
+By construction, Stein’s shrinkage improves Lipschitzness and therefore robustness of the model, while maintaining high accuracy.
+Acknowledgments and Disclosure of Funding
+This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099 to Vemuri.
+It was also partially supported by the NSF Smart and Connected Health grant 2205417 to Fletcher.
+PPMI Data Availability Statement: Data used in the preparation of this article were obtained from the Parkinson’s Progression Markers Initiative (PPMI) database [(https://www.ppmi-info.org/access-dataspecimens/download-data, accessed on 6 December 2023, RRID:SCR 006431].
+For up-to-date information on the study, visit www.ppmi-info.org.
+This analysis used data openly available from PPMI, obtained from PPMI upon request.
+PPMI Data Acknowledgments: We extend our gratitude to Lesa Melnyczuk and Iryna Milne for their valuable assistance.
+Data used in the preparation of this article were obtained on 6 December, 2023 from the Parkinson’s Progression Markers Initiative (PPMI) database (https://www.ppmi-info.org/access-data-specimens/download-data, accessed on 6 Decemeber 2023), RRID:SCR-006431.
+For up-to-date information on the study, visit http://www.ppmi-info.org, accessed on 6 Dec.
+2023.
+PPMI—a public–private partnership—is funded by the Michael J.
+Fox Foundation for Parkinson’s Research and funding partners, including 4D Pharma, Abbvie, AcureX, Allergan, Amathus Therapeutics, Aligning Science Across Parkinson’s, AskBio, Avid Radiopharmaceuticals, BIAL, BioArctic, Biogen, Biohaven, BioLegend, BlueRock Therapeutics, Bristol-Myers Squibb, Calico Labs, Capsida Biotherapeutics, Celgene, Cerevel Therapeutics, Coave Therapeutics, DaCapo Brainscience, Denali, Edmond J.
+Safra Foundation, Eli Lilly, Gain Therapeutics, GE HealthCare, Genentech, GSK, Golub Capital, Handl Therapeutics, Insitro, Jazz Pharmaceuticals, Johnson & Johnson Innovative Medicine, Lundbeck, Merck, Meso Scale Discovery, Mission Therapeutics, Neurocrine Biosciences, Neuron23, Neuropore, Pfizer, Piramal, Prevail Therapeutics, Roche, Sanofi, Servier, Sun Pharma Advanced Research Company, Takeda, Teva, UCB, Vanqua Bio, Verily, Voyager Therapeutics, the Weston Family Foundation, and Yumanity Therapeutics.
 Appendix A Technical Appendices and Supplementary Material
-In this Appendix, we present the detailed proofs of the two key theorems 3.1 and 3.3 and additional experiments showing Stein’s correction applied only to the mean and not the variance.
-These experimental results are depicted here in this Appendix as additional rows added in to the three tables appearing in the main body of the paper.
-Also included are further experimental details.
-Theorem 3.1 Proof:.
-We denote MLE estimator as $\hat{\theta}^{\mathrm{MLE}}=z=\left(z_{1},\ldots z_{p}\right).$
-Then James - Stein Estimator is $\hat{\theta}^{\mathrm{JS}}=z-\frac{(p-2)z}{z^{\top}z}=z\left(1-\frac{p-2}{z^{\top}z}\right)=\hat{\theta}^{\mathrm{MLE}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{MLE}\top}\hat{\theta}^{\mathrm{MLE}}}\right).$
-The risk of James - Stein Estimator equals to
-$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$
+In this Appendix, we present the detailed proofs of the two key theorems 1 and 3, and provide a detailed discussion of the experimental setups.
+Proof [Theorem 1 Proof:]
+We denote the observation-based estimator by
+$\hat{\theta}^{\mathrm{0}}=Z=(Z_{1},\ldots,Z_{p}).$
+Then the James–Stein estimator is
+$\hat{\theta}^{\mathrm{JS}}=Z-\frac{(p-2)Z}{Z^{\top}Z}=Z\left(1-\frac{p-2}{Z^{\top}Z}\right)=\hat{\theta}^{\mathrm{0}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{0}\top}\hat{\theta}^{\mathrm{0}}}\right).$
+The risk of James - Stein Estimator is given by,
+$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$ (7)
 $=\mathbb{E}\left(\left\{(Z-\theta)-\frac{(p-2)Z}{Z^{\top}Z}\right\}^{\top}\left\{(Z-\theta)-\frac{(p-2)Z}{Z^{\top}Z}\right\}\right)=$
 $=\mathbb{E}\left((Z-\theta)^{\top}(Z-\theta)\right)-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}{\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)}=$
-$=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)$
+$=R(\hat{\theta}^{\mathrm{0}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)$
 Let us consider the last two terms:
-$-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
+$-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=$ (8)
 $=-2(p-2)\left(1-\mathbb{E}\left(\frac{Z^{\top}\theta}{Z^{\top}Z}\right)\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=$
 $=(p-2)\left(-2+\mathbb{E}\left(\frac{2Z^{\top}\theta}{Z^{\top}Z}\right)+\mathbb{E}\left(\frac{(p-2)}{Z^{\top}Z}\right)\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\right)$
-We want to show that $\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)<2$
-Let us denote $\tilde{x}=2z^{\top}\theta+p-2\quad$ and $\tilde{y}=\frac{1}{z^{\top}z}.$
-By applying Cauchy–Schwarz inequality we get,
-$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)=\mathbb{E}(\tilde{x}\cdot\tilde{y})\leq|\mathbb{E}(\tilde{x}\cdot\tilde{y})|\leq\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}.$
-Let us calculate each term separately:
-$\mathbb{E}\left(\tilde{x}^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta+p-2\right)^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta\right)^{2}+4\left(z^{\top}\theta\right)(p-2)+(p-2)^{2}\right)=$
-$=\mathbb{E}\left(4\left(x^{\top}\theta+y^{\top}\theta\right)^{2}+4\left(x^{\top}\theta+y^{\top}\theta\right)(p-2)\right)+(p-2)^{2}=$
-$=4\mathbb{E}\left((x^{\top}\theta)^{2}\right)+8\mathbb{E}\left((x^{\top}\theta)(y^{\top}\theta)\right)+4\mathbb{E}\left((y^{\top}\theta)^{2}\right)+4(p-2)\mathbb{E}\left(x^{\top}\theta\right)+(p-2)^{2}=$
-$=4Var\left(x^{\top}\theta\right)+4\left(\mathbb{E}\left(x^{\top}\theta\right)\right)^{2}+4Var\left(y^{\top}\theta\right)+4\left(\mathbb{E}\left(y^{\top}\theta\right)\right)^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=$
-$=4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+4Var\left(y^{\top}\theta\right)+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}\leq$
-$4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+8\epsilon^{2}||\theta||_{2}^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}.$
-So, $\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\leq\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}.$
-Note that,
-$\mathbb{E}\left(z^{\top}z\right)=\mathbb{E}\left(\sum_{i}z_{i}^{2}\right)=\sum_{i}\mathbb{E}\left(z_{i}^{2}\right)=\sum_{i}Var(z_{i})+\sum_{i}\left(\mathbb{E}(z_{i})\right)^{2}\geq p+||\theta||_{2}^{2}.$
-Using Jensen inequality, firstly for $f(x)=\frac{1}{x}$ and then for $g(x)=x^{2}$ we get
-$\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}=\sqrt{\mathbb{E}\left(\frac{1}{(z^{\top}z)^{2}}\right)}\leq\sqrt{\frac{1}{\mathbb{E}\left((z^{\top}z)^{2}\right)}}\leq\frac{1}{\mathbb{E}\left(z^{\top}z\right)}\leq\frac{1}{p+||\theta||_{2}^{2}}.$
-So, overall we get
-$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\leq\frac{\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}}{p+||\theta||_{2}^{2}}\leq 2.$
-The final inequality can be established by direct computation. ∎
-Theorem 3.3 Proof:.
+We want to show that
+$\mathbb{E}\!\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$
+For convenience, set
+$\tilde{x}=2Z^{\top}\theta+p-2,\quad\tilde{y}=\frac{1}{Z^{\top}Z}.$
+Then,
+[…]
+By the Cauchy–Schwarz inequality,
+$\mathbb{E}(\tilde{x}\,\tilde{y})\leq\sqrt{\mathbb{E}(\tilde{x}^{2})}\,\sqrt{\mathbb{E}(\tilde{y}^{2})}.$
+Let us calculate each term separately. First,
+$\mathbb{E}\!\left(\tilde{x}^{2}\right)$ $=\mathbb{E}\!\left(\big(2Z^{\top}\theta+p-2\big)^{2}\right)$ (9)
+$=\mathbb{E}\!\left((2Z^{\top}\theta)^{2}+4(p-2)(Z^{\top}\theta)+(p-2)^{2}\right)$
+$=4\,\mathbb{E}\!\left((Z^{\top}\theta)^{2}\right)+4(p-2)\,\mathbb{E}(Z^{\top}\theta)+(p-2)^{2}.$
+Recall that $Z=X+Y$, where $X\sim\mathcal{N}(\theta,I_{p})$ with mean $\theta$ and $Y$ is a mean-zero sub-Gaussian noise. Hence
+$\mathbb{E}\!\left((Z^{\top}\theta)^{2}\right)$ $=\mathbb{E}\!\left((X^{\top}\theta+Y^{\top}\theta)^{2}\right)$ (10)
+$=\mathbb{E}\!\left((X^{\top}\theta)^{2}\right)+2\,\mathbb{E}\!\left((X^{\top}\theta)(Y^{\top}\theta)\right)+\mathbb{E}\!\left((Y^{\top}\theta)^{2}\right).$
+Now
+$\mathbb{E}\!\left((X^{\top}\theta)^{2}\right)=\operatorname{Var}(X^{\top}\theta)+\big(\mathbb{E}(X^{\top}\theta)\big)^{2}=\|\theta\|_{2}^{2}+\|\theta\|_{2}^{4},$ (11)
+and
+$\mathbb{E}(X^{\top}\theta)=\|\theta\|_{2}^{2}.$ (12)
+For the noise term, by sub-Gaussianity we get,
+$\mathbb{E}\!\left((Y^{\top}\theta)^{2}\right)\leq 2\varepsilon^{2}\,\|\theta\|_{2}^{2}.$ (13)
+Substituting (10)–(13) into (9) we obtain,
+$\mathbb{E}\!\left(\tilde{x}^{2}\right)\leq 4\|\theta\|_{2}^{2}\left(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\right)+(p-2)^{2}.$
+Therefore,
+$\sqrt{\mathbb{E}\!\left(\tilde{x}^{2}\right)}\;\leq\;\sqrt{\,4\|\theta\|_{2}^{2}\big(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\big)+(p-2)^{2}}.$ (14)
+First, observe that
+$\mathbb{E}\!\left(Z^{\top}Z\right)=\sum_{i=1}^{p}\mathbb{E}(Z_{i}^{2})=\sum_{i=1}^{p}\operatorname{Var}(Z_{i})+\sum_{i=1}^{p}\big(\mathbb{E}(Z_{i})\big)^{2}\;\;\geq\;p+\|\theta\|_{2}^{2}.$
+Next, by Jensen’s inequality applied to the convex functions $f(x)=1/x$ and $g(x)=x^{2}$, we have
+$\sqrt{\mathbb{E}\!\left(\tilde{y}^{2}\right)}$ $=\sqrt{\mathbb{E}\!\left(\frac{1}{(Z^{\top}Z)^{2}}\right)}$ (15)
+$\leq\sqrt{\frac{1}{\mathbb{E}\!\left((Z^{\top}Z)^{2}\right)}}\;\leq\;\frac{1}{\mathbb{E}(Z^{\top}Z)}$
+$\leq\frac{1}{p+\|\theta\|_{2}^{2}}.$
+Combining (14) and (15), we obtain
+$\mathbb{E}\!\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\;\leq\;\frac{\sqrt{\,4\|\theta\|_{2}^{2}\big(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\big)+(p-2)^{2}}}{p+\|\theta\|_{2}^{2}}\;\leq\;2.$
+The last inequality can be verified by direct computation.
+Proof [Theorem 3 Proof:]
 […]
 Note that,
-[…]
-[…]
+$R(\hat{\beta}^{\,\mathrm{JS}},\beta)=\mathbb{E}\left(\left\|\hat{\beta}^{\,\mathrm{JS}}-\beta\right\|_{2}^{2}\right)=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}+cV-\beta\right\|_{2}^{2}\right)=$
+$=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}-\beta\right\|_{2}^{2}\right)+2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)\cdot cV\right)+p\mathbb{E}\left(c^{2}V^{2}\right)$
 So, we need to show:
@@
     $c<0$ and
-    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (7)
+    $2\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>0$ (16)
   • Case 2:
     $c>0$ and
-    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (8)
-Let us derive some facts that will be helpful in the proof.
-Firstly, we aim to show that for each $i$, the following inequality holds:
+    $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<0$ (17)
+We now establish an intermediate result in the form of a lemma that will be useful in the proof of Theorem 3.
+Lemma 6 (Probability bound for $\hat{\sigma}^{2}_{y,i}+W_{i}$)
+Let $X_{ij}\sim\mathcal{N}(\mu,\sigma_{x,i}^{2})$ and $Y_{ij}\sim\mathrm{SG}(2\varepsilon^{2})$ be independent for $j=1,\dots,n$. Define
+$\hat{\sigma}_{y,i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(Y_{ij}-\bar{Y}_{i})^{2},\quad W_{i}=\frac{2}{n}\sum_{j=1}^{n}(X_{ij}-\bar{X}_{i})(Y_{ij}-\bar{Y}_{i}),$
+where $\bar{X}_{i}=\frac{1}{n}\sum_{j}X_{ij}$, $\bar{Y}_{i}=\frac{1}{n}\sum_{j}Y_{ij}$. Then, for any $\delta\in(0,1)$,
 $\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta.$
+where $\delta\to 0$ as $n\to\infty$.
+Proof
 Note that:
@@
 $\nu^{2}=C_{1}\cdot\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2},\quad b=C_{2}\cdot\sqrt{\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2}}.$
-Reminder. A random variable $Z$ is sub-exponential if:
-$\mathbb{E}[e^{\lambda Z}]\leq\exp\left(\frac{\lambda^{2}\nu^{2}}{2}\right),\quad|\lambda|<\frac{1}{b}.$
-Bernstein Inequality for Sub-Exponential Random Variables. Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential variables with parameters $(\nu,b)$. Then for any $t>0$:
+We now recall the Bernstein inequality that will be used to prove the desired bound.
+Bernstein Inequality for Sub-Exponential Random Variables: Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential [^1] A random variable $Z$ is sub-exponential with parameters $(\nu,b)$ if $\mathbb{E}[e^{\lambda Z}]\leq\exp(\lambda^{2}\nu^{2}/2)$ for all $|\lambda|<1/b$.
+variables with parameters $(\nu,b)$.
+Then for any $t>0$:
 $\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\leq-t\right)\leq\begin{cases}\exp\!\left(-\dfrac{nt^{2}}{2\nu^{2}}\right),&\text{if }0\leq t\leq\dfrac{\nu^{2}}{b},\\ \exp\!\left(-\dfrac{nt}{2b}\right),&\text{if }t>\dfrac{\nu^{2}}{b}.\end{cases}$
@@
 $\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta,$
-where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$.
-Thus, the desired non-negativity condition holds with high probability in the large sample regime.
-We will use this fact in the proof of case 2.
-Now we are ready to consider each case separately.
-Case 1. We examine the first term in more detail:
+where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$. Thus, the desired non-negativity condition holds with high probability in the large sample regime.
+We will use this lemma in the proof of case 2. Now we are ready to consider each case separately.
+Case 1. We examine the first term of (16) in more detail:
 $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]=2\sum_{i=1}^{p}\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}-\beta_{i}\right)V\right]$ (18)
 $=2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)$
-Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can apply Stein’s lemma for the Gamma distribution:
-Lemma A.1 (Stein’s Lemma for the Gamma Distribution).
-Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that
-$\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty.$
-Then the following identity holds:
+Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can now apply Stein’s Lemma for the Gamma Distribution, which states: Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty$ and $\mathbb{E}[Xh^{\prime}(X)]<\infty.$ Then the following identity holds:
 $\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)].$
-We apply Stein’s identity to the function $h(x)=V$, where
-$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}.$
-Using the Gamma Stein identity:
-$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$
-we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ as follows:
+This can be rewritten as:
+$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)].$ (19)
+We now apply Stein’s identity in (19) to the function $h(X)=V$, where
+$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)\right]^{1/p}.$
+The derivative of $V$ with respect to $\hat{\sigma}_{x,k}^{2}$ is:
 $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+W_{k}}.$
-Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
-$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
-In particular, we obtain the inequality:
-$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$
-Using this bound equation (9) becomes:
+Substituting Stein’s identity (19) for $\hat{\sigma}_{x,i}^{2}$, we get the following.
+$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}\right].$ (20)
+In particular, we can obtain the following inequality:
+$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$ (21)
+Using the above bound in (21), equation (18) becomes:
 $2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)\geq$
 $2\sum_{i=1}^{p}\left(\mathbb{E}\left[\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}V\right]-\beta_{i}\cdot\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)=$
-$=2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
-$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$
-Let’s denote
-$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$
-Overall, we get lower bound for :
+$2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
+$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$ (22)
+Recall that in the statement of the theorem we introduced $\hat{\sigma}_{z,i}^{2}=\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}$.
+For convenience, in the proof we will work with the averaged version
+$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$ (23)
+From (22) the lower bound for (16) is given by:
 $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\mathbb{E}[V^{2}]\geq-\frac{2p}{\alpha(\alpha+1)}\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]+pc\mathbb{E}[V^{2}]\geq$
 $\geq-p\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]\left(\frac{2}{\alpha(\alpha+1)}+c\right)>0$
-Last inequality appears due to assumption. So we have:
-$c<0\ \text{and}\ \frac{2}{\alpha(\alpha+1)}+c<0\Rightarrow c\leq-\frac{2}{\alpha(\alpha+1)}.$
+The last inequality follows from our case-1 hypothesis. Therefore,
+[…]
 Case 2. We assume $c>0$ and we want to show
-$02\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<$ 0 .
-Recall that using Bernstein’s inequality we get with high probability, the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative.
-So, combining this with Stein’s lemma for $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ we get:
+$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<0.$
+Recall that using the lemma 6, with high probability, we find that the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative.
+So, combining this with Stein’s lemma (19) applied to $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ and $V$ we get:
 $\mathbb{E}(V)$ $=\frac{1}{\alpha\beta_{i}}\,\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\,\mathbb{E}\!\left(V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}\right)$
-$\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}(V)$
-$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$
+$\geq\frac{1}{\alpha\beta_{i}}\,\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\,\mathbb{E}(V).$
+$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$ (24)
 $\Rightarrow\quad-\mathbb{E}(V)$ $\leq-\frac{1}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\cdot\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V).$
-Plug in this bound and get inequality for  (9):
+Substituting the bound (24) into (18), we get the following inequality:
 $2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)$ $=2\sum_{i=1}^{p}\mathbb{E}\left(\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right)\right)\right)$
-$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$
-Using our previous notation $\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2},$
-$=\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$
-Overall for (8) we get:
+$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$ (25)
+As defined in (23) $\overline{\hat{\sigma}_{z}^{2}}:=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}$ , we can rewrite the expression (25) as
+$\frac{2p}{\alpha+1}\,\mathbb{E}\!\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha\left(1+\frac{1}{\alpha p}\right)}\,\mathbb{E}\!\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$ (26)
+So, for (17) we get the following bound by applying (26):
 $2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})\leq$
 $\leq\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)+pc\,\mathbb{E}(V^{2})<0$
-The last inequality we want to establish by finding the appropriate $c$. So, we will rewrite it in a more suitable way:
+By finding the appropriate $c$ we can obtain the last inequality.
+Rewriting it in a more cleaner way, we have:
 $c\,\mathbb{E}(V^{2})<\frac{2}{\alpha p+1}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)-\frac{2}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)$
 $c<\frac{2}{\alpha p+1}\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}-\frac{2}{\alpha+1}\cdot\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}$ (27)
-Let’s bound each term independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, then we bound second term
+Let us now bound each term (on the RHS of the above inequality) independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, so we can write a bound for second term in (27)
 $\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)\geq\mathbb{E}(V^{2})\quad\Rightarrow\quad-\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}\leq-1.$ (28)
 Also, using Cauchy–Schwarz inequality:
-$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$
-For the first term we get:
+$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$ (29)
+Using (29) we can bound the first term in (27):
 $\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}\cdot\sqrt{\mathbb{E}(V^{2})}}{\mathbb{E}(V^{2})}=\sqrt{\frac{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}{\mathbb{E}(V^{2})}}$ (30)
-By applying Jensen’s inequality for $f(x)=\sqrt{x}$ – concave, we have:
+By applying Jensen’s inequality to the function of the form: $f(x)=\sqrt{x}$, we have:
 $\sqrt{\mathbb{E}(V^{2})}\geq\mathbb{E}(|V|)=\mathbb{E}(V)\Rightarrow\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}$ (31)
-Now,
-$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq$
-Again, using Jensen’s inequality for $f(x)=\exp{x}$
-$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
-Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability
-$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$
-We know that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
-$\mathbb{E}(\log\sigma_{x,i}^{2})\ \geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$
-Plug in this inequality to the derivations above we get:
-$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$
-For (13) we get:
-$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$
-Also, we can directly calculate numerator of (12):
+Again, applying Jensen’s inequality to the function $f(x)=\exp{x}$ in the following expression, we get,
+$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
+Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability (see lemma 6) we get:
+$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$ (32)
+Recall that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is a digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can use the bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
+$\mathbb{E}(\log\sigma_{x,i}^{2})\geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$ (33)
+Substituting the bound (33) into (32) we obtain the following inequality:
+$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$ (34)
+Now, using (34) for (31) we get:
+$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$ (35)
+Also, we can directly calculate the numerator of (30):
 Firstly,
@@
 $\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2})=\alpha^{2}\beta_{i}\beta_{j}.$
+Then the numerator of (30) is given by:
 $\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}=\sqrt{\sum_{i=1}^{p}\mathbb{E}(\hat{\sigma}_{x,i}^{4})+2\sum_{1\leq i<j\leq p}\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2}})=$
 $=\sqrt{\sum_{i=1}^{p}\alpha\beta_{i}^{2}(1+\alpha)+2\sum_{1\leq i<j\leq p}\alpha^{2}\beta_{i}\beta_{j}}=\sqrt{\alpha(1+\alpha)\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}\left((\sum_{i=1}^{p}\beta_{i})^{2}-\sum_{i=1}^{p}\beta_{i}^{2}\right)}=$
-$=\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$
-Therefor, for (12) we get upper bound:
-$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$
-Using this inequality and (11) finally for (10) we get:
+$\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$ (36)
+Therefore, combining (36) and (35) we get an upper bound for (30):
+$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$ (37)
+Finally, for (27) using inequality (37) and (28) we get:
 $c<\frac{2}{\alpha p+1}\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}-\frac{2}{\alpha+1}=$
-$=\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$
+$\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$ (38)
 Note, that
@@
 $\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\beta_{i}}\leq\frac{1}{\frac{1}{p}\sum_{i=1}^{p}\beta_{i}}=\frac{p}{\sum_{i=1}^{p}\beta_{i}}.$
-Hence,
+Hence, we can write a bound for the first term in (38):
 $\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\leq\frac{1}{p}\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}\left(\sum_{i=1}^{p}\frac{1}{\beta_{i}}\right)$
-$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$
-So, dominance holds for
+$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$ (39)
+According to (38) in combination with (39) dominance holds for
 $0<c<\frac{2p}{(\alpha p+1)}\exp{\left(\frac{1}{\alpha}\right)}\sqrt{\frac{1}{\alpha}+1}-\frac{2}{\alpha+1}.$
-Note that under regular conditions Stein’s estimate dominance holds when constant $c$ satisfies
+Note that under regular conditions the dominance of Stein’s estimate holds when constant $c$ satisfies
 $0<c<\frac{2p}{\alpha p+1}-\frac{2}{\alpha+1}=\frac{2(p-1)}{(\alpha+1)(\alpha p+1)}.$
@@
 Appendix B Experiments
-We will, firstly, provide derivations of Lasso and Ridge shrinkage estimators for mean and variance.
-Lasso mean shrinkage. We consider the following penalized least squares problem:
-$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda|\mu|\right),$
-where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage.
-This is equivalent to applying an $l_{1}$ – penalty to the mean and the solution is given by the soft-thresholding operator:
-$\hat{\mu}=sign(\bar{x})\max{(0,|\bar{x}|-\frac{\lambda}{2n}),}$
-where $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ is the sample mean. This form encourages shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large.
-Lasso variance shrinkage. A similar penalized formulation can be used for estimating the variance, leading to a soft-thresholding-type estimator for sample variance $s^{2}$:
-$\hat{\sigma}^{2}=\max{(0,s^{2}-\frac{\lambda}{2}),}$
-where $\lambda>0$ is regularization hyperparameter, $s^{2}=\frac{1}{n}\sum_{i}(x_{i}-\bar{x})^{2}.$
-Ridge mean shrinkage. We consider the following penalized least squares problem:
-$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda\mu^{2}\right),$
-where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage.
-This is equivalent to applying an $l_{2}$ – penalty to the mean.
-Taking the derivative and setting it to zero yields the following solution:
-$\hat{\mu}=\frac{\sum_{i}x_{i}}{n+\lambda}.$
-Ridge variance shrinkage. Similarly, we can obtain the ridge estimate for the variance:
-$\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$
-where $s^{2}$ is the sample variance and $\lambda>0$ is the regularization parameter.
-Now let us discuss reproducibility of experiments on each dataset.
-B.1 CIFAR10 dataset
-We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes.
-We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing.
-Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 4.
-We used the standard ResNet-9 architecture, modifying only the batch normalization layers according to the method under consideration.
-For data splits, we used the standard predefined training and test sets.
-Training time was approximately $10$ minutes per run, with negligible variation across methods.
-All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM.
-Each experiment was repeated $11$ times to estimate the standard deviation of the performance metrics, and we tested three different batch sizes $(32,64,128)$, resulting in a total training time of approximately $1800$ minutes.
-We used SGD with Nesterov acceleration and default parameters, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$.
-We employed early stopping based on validation accuracy on data without noise to prevent overfitting.
-We used cross-entropy loss for training, and report classification accuracy as the main evaluation metric.
-Finally, for the Lasso and Ridge shrinkage methods, we selected the regularization parameter $\lambda\in(10^{-3},10^{-1})$ yielding the best performance on validation dataset without noise.
-Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
-Standard BN (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
-Standard BN (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
-Standard BN (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
-Our BN (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
-Our BN (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
-Our BN (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
-Mean corrected BN (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
-Mean corrected BN (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
-Mean corrected BN (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
-Khoshsirat et al. BN [15] (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
-Khoshsirat et al. BN [15] (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
-Khoshsirat et al. BN [15] (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
-Lasso corrected BN (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
-Lasso corrected BN (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
-Lasso corrected BN (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
-Ridge corrected BN(32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
-Ridge corrected BN(64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
-Ridge corrected BN(128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
-Table 4: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
-B.2 Cityscapes dataset
-We performed segmentaion experiments on Cityscapes dataset – high-resolution urban scene segmentation with fine-grained object boundaries.
-The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation.
-In all our experiments we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics.
-Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 5.
-We use the HRNetV2-W16 architecture (where 16 denotes the width of the high-resolution convolution) as our baseline model.
-The original training set was randomly split into $80\%$ for training and $20\%$ for validation, while the original validation set was used as the test set.
-All images were resized from $1024\times 2048$ to $512\times 1024$ for computational efficiency.
-To account for variability and estimate the standard deviation of our results, each experiment was repeated $7$ times for each batch size $(16,32).$ Each run took approximately $1.5$ hours on a single NVIDIA A100 GPU with $64$ GB of RAM and $8$ CPU cores, resulting in a total training time of around 126 hours.
-We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$.
-Early stopping based on the mean Intersection over Union (mIoU) score on validation data without noise was employed to prevent overfitting.
-We used cross-entropy loss for training and report mIoU as the primary evaluation metric.
-Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise.
-Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
-Standard BN (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
-Standard BN (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
-Our BN (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
-Our BN (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
-Mean corrected BN (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
-Mean corrected BN (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
-Khoshsirat et al. BN [15] (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
-Khoshsirat et al. BN [15] (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
-Lasso corrected BN (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
-Lasso corrected BN (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
-Ridge corrected BN(8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
-Ridge corrected BN(16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
-Table 5: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
-B.3 Details of Application to PPMI Data
-In this experiment, the goal is to demonstrate the performance of our BN with Stein shrinkage on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls).
-This data was obtained from the PPMI (Parkinson’s Progressive Markers Initiative) portal after a formal request was filled out.
-Please see data access citation and acknowledgment below for this data.
-DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged.
-The PPMI data was subjected to a standard pre-processing pipeline involving Eddy currents and motion correction etc., using the FSL software [13].
-DTIs are matrix-valued images in which each voxel takes on a value that is a $(3,3)$ symmetric positive definite (SPD) matrix representing the covariance matrix of a zero-mean local Gaussian at the voxel.
-These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps.
-The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI.
-Harmonization is applied to these scalar feature maps to remove site effects caused by distinct acquisition protocols and equipment used.
-Harmonization was achieved using the method called ComBat adapted to diffusion tensor images and associated software presented in [12].
-It should be noted that there are other sophisticated hamronization methods that one can apply to the data but this one is simple.
-The data set is imbalanced and consists of $490$ (130 PD and 360 Controls) FA maps.
-To address this, we use mixup-based[31] augmentation with sampling bias favoring the minority class (PD), this approach allowed us to generate additional training examples.
-We follow the same procedure in our other experiments for the setup and use a 5-fold cross validation to compare 3D CNN models with different variants of 3D BatchNorm layer.
-The architecture consists of four 3D convolutional layers, each followed by a 3D BatchNorm layer, and three linear layers for binary classification.
-The 3D input tensors have spatial dimensions of $(128\times 128)$ and a depth of $64$.
-We conduct experiments using two batch sizes ($10$ and $16$) and optimize the binary cross-entropy loss.
-Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 6.
-We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$.
-Early stopping based on the accuracy score on validation data without noise was employed to prevent overfitting.
-Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise.
-All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM.
-Each configuration take around $20$ minutes, which end up to the 20 hours for the whole experiments.
-Method 3% noise 10% noise 15% noise 20% noise 30% noise
-Standard BN (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
-Standard BN (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
-Our BN (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
-Our BN (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
-Mean corrected BN (10) 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
-Mean corrected BN (16) 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
-Khoshsirat et al. BN [15] (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
-Khoshsirat et al. BN [15] (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
-Lasso corrected BN (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
-Lasso corrected BN (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
-Ridge corrected BN(10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
-Ridge corrected BN(16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
-Table 6: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data. Values depict mean of accuracies $\pm$ standard deviation.
+Unless otherwise specified, all models were trained with stochastic gradient descent (SGD) with Nesterov acceleration, cross-entropy loss, and early stopping based on validation performance on noise-free data.
+The learning rate was adjusted separately for each method in the range $(10^{-5},10^{-2})$.
+For the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on the noise-free validation performance.
+All experiments were performed on NVIDIA A100 GPUs with CPU and RAM resources adjusted to the size of the data set.
+To assess variability, each setting was repeated multiple times and we report averages with standard deviations.
+B.1 CIFAR-10
+[…]
+B.2 Cityscapes
+We used the HRNetV2-W16 architecture with images resized from $1024\times 2048$ to $512\times 1024$.
+The dataset contains 5,000 finely annotated images across 19 classes.
+We split the original training set into $80\%$ training and $20\%$ validation, and used the original validation set as test data.
+Sub-Gaussian noise was injected into feature maps during validation and testing.
+We tested batch sizes $(16,32)$, with each experiment repeated $7$ times.
+Each run took $\sim$1.5 hours, totaling $\sim$126 GPU hours.
+B.3 PPMI Data
+We employed 3D CNN architecture consisting of four convolutional layers, each followed by a 3D BN variant under consideration, and three fully connected layers for binary classification.
+The input tensors had spatial dimensions of $(128\times 128)$ and a depth of $64$.
+We tested batch sizes of $10$ and $16$, using 5-fold cross-validation.
+Each configuration took $\sim$20 minutes, for a total of $\sim$20 GPU hours.
+B.4 Comparison with data dependent attacks
+For data dependent adversarial attack experiments, we use trained models with their best-performing hyperparameters and apply attacks only to the test data.
+The overall computational setup, including runtime, memory requirements, and hardware, remains the same as in the previous experiments.
+As mentioned earlier, both FGSM and PGD attacks are evaluated at different perturbation levels whose values range over, ($\epsilon=0.03,0.1,0.15,0.2,0.3$).
+For PGD, we additionally specify attack hyperparameters: on the CIFAR-10 dataset, we use 5 iterations with a fixed step size $\alpha=1/255$, while on the Cityscapes dataset we use 5 iterations with an adaptive step size $\alpha=\epsilon/T$, where $T$ is the total number of iterations.
+The corresponding results are reported in Tables 4, 5, 6, and 7 respectively.